### v2

# HTMformer: Hybrid Time and Multivariate Transformer for Time Series Forecasting

Abstract
Transformer-based methods have achieved impressive results in time series forecasting. However, existing Transformers still exhibit limitations in sequence modeling as they tend to overemphasize temporal dependencies. This incurs additional computational overhead without yielding corresponding performance gains. We find that the performance of Transformers is highly dependent on the embedding method used to learn effective representations. To address this issue, we extract multivariate features to augment the effective information captured in the embedding layer, yielding multidimensional embeddings that convey richer and more meaningful sequence representations. These representations enable Transformer-based forecasters to better understand the series. Specifically, we introduce Hybrid Temporal and Multivariate Embeddings (HTME). The HTME extractor integrates a lightweight temporal feature extraction module with a carefully designed multivariate feature extraction module to provide complementary features, thereby achieving a balance between model complexity and performance. By combining HTME with the Transformer architecture, we present HTMformer, leveraging the enhanced feature extraction capability of the HTME extractor to build a lightweight forecaster. Experiments conducted on eight real-world datasets demonstrate that our approach outperforms existing baselines in both accuracy and efficiency.
1 Introduction
Figure 1: Performance (MSE) of HTMformer. Smaller image area yields better performance.
Long-term time series forecasting holds significant importance in various fields such as finance and economics (Andersen et al., 2006), climate science (Mudelsee, 2019), healthcare (Zeger et al., 2006), geophysics (Gubbins, 2004), industrial monitoring (Truong et al., 2022), etc. Recently, Transformer (Vaswani et al., 2017) and its variants have achieved tremendous success in time series forecasting, showing great promise. Current Transformer-based forecasters often overemphasize modeling temporal dependencies. However, the effective information within the temporal dimension is limited. Consequently, existing models incur increasing computational overhead without significant accuracy improvements.
Unlike natural language sequences, time series data, particularly in domains such as transportation, meteorology, and electricity systems, are influenced by specific physics-based systems, e.g., traffic flow is largely governed by the real-world road network. Therefore, multivariate correlations are essential for time series forecasting tasks (Zheng & Sun, 2024). Two key points are worth noting: 1) the vast majority of features in time series are contained within the temporal dimension, and the depth of exploration in this dimension significantly influences prediction accuracy (Hyndman & Athanasopoulos, 2021); 2) the multivariate dimension also contains a large number of features (Pravilovic et al., 2013), and fully exploiting multivariate features can further boost performance.
The performance of Transformer-based forecasters is highly dependent on the design of the embedding layer. Before applying multi-head self-attention to the embeddings, effective feature extraction is critical to enhancing the model’s sequence modeling capacity (Li et al., 2023a). We compare the ground-truth values and the predictions obtained under different embedding strategies on four datasets (see Appendix A). Specifically, models such as iTransformer (Liu et al., 2024b) and PatchTST (Nie et al., 2023), which employ an inverted embedding strategy or patch-based embedding strategy for deeper time-series representation learning, achieve significantly better performance than the vanilla Transformer, which only uses simple convolutional embeddings. Meanwhile, we also notice that as the input length increases, the prediction loss fails to markedly decrease and even tends to rise. These suggest that the modeling capacity of Transformer-based forecasters is inherently constrained by the quality of embeddings, and this limitation cannot be eliminated by increasing the lookback length.
To mitigate this issue, this paper proposes to jointly capture temporal and multivariate features in the embedding layer, thereby yielding semantically rich embedding representations to overcome the inherent limitations of the conventional embedding strategy (see Appendix B). Utilizing proposed strategy, we constructed a Transformer-based forecaster and achieved state-of-the-art empirical performance on benchmarks, as shown in Figure 1. The contributions of this work are as follows:
  • We propose the Hybrid Temporal and Multivariate Embedding (HTME) extractor, which comprises a lightweight temporal feature extractor to capture informative temporal dynamics and a multivariate feature extractor to model the multivariate dependencies. The HTME is broadly applicable across Transformer-based forecasting architectures and effectively enhances the capacity of transformer-based models to model complex time series.
  • We introduce a novel forecasting model, HTMformer, to further validate the effectiveness of the HTME strategy. In this framework, the HTME serves as input to a vanilla encoder-only Transformer framework. Extensive empirical evaluations are conducted on eight real-world benchmarks spanning application domains. Notably, HTMformer consistently achieves state-of-the-art performance across various benchmarks.
2 related Work
Transformer-Based Time Series Forecasting.
Well-designed Transformer-based forecasters have achieved substantial success in time series forecasting tasks. They are not limited by the inability of CNN-based methods (Liu et al., 2022) to capture long-term temporal dependencies (Sutskever et al., 2014; Karim et al., 2017), nor do they encounter the vanishing and exploding gradient problems (Bengio et al., 1994; Hochreiter & Schmidhuber, 2019) commonly observed in RNN-based methods (Cho et al., 2014; Lai et al., 2018; Salinas et al., 2020). In recent years, advanced models such as Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), and Crossformer (Zhang & Yan, 2023) have primarily introduced novel attention mechanisms and architectural modifications for explicitly modeling long-term dependencies and temporal interactions in time series data. Nonetheless, as discussed in the Introduction, Transformer-based forecasters suffer from inherent limitations in effectively modeling time series. Refining the attention mechanism and architecture alone cannot address this limitation. Unlike prior works, we propose a novel embedding strategy for comprehensive feature extraction on the raw input before it is fed into the attention module, thereby substantially improving the modeling capacity of diverse attention mechanisms. As shown in Figure 2.
Temporal Feature Extraction.
Previous Transformer-based forecasters, such as FEDformer (Zhou et al., 2022) and MultiPatchFormer (Naghashi et al., 2025), tend to overprioritize temporal feature extraction, introducing various complex mechanisms to model temporal dependencies. This design often incurs substantial computational overhead without delivering notable performance gains. We argue that the informational content in the temporal dimension is inherently limited and should not be excessively emphasized. In contrast, HTMformer performs full temporal feature extraction solely within the embedding layer, thereby significantly reducing computational complexity.
Multivariate Correlation.
iTransformer retains all native Transformer components without modification. Instead, it adopts an inverted input representation, enabling the self-attention model to better capture multivariate correlations and thereby yielding satisfactory performance. This underscores the importance of modeling multivariate correlations in capturing the semantics of time series. In CNN-based predictors, graph convolutional networks (Kipf & Welling, 2017; Yu et al., 2018; Guo et al., 2019) explicitly model correlations among variables by leveraging graph adjacency matrices. This paradigm unifies the temporal dynamics of the data with the multivariate structure of variables within a graph-based model, thereby substantially improving performance in time series tasks. However, such explicit modeling entails significant computational overhead, particularly in Transformer architectures. HTMformer is designed to moderately extract multivariate features and integrate them with temporal information within the embedding layer, yielding semantically enriched embeddings. Furthermore, it employs an inverted input strategy to enhance the efficiency and performance of Transformers and reduce the computational costs.
Figure 2: Previous works have focused on redesigning Transformer architectures or attention mechanisms. We propose a novel embedding strategy.
3 HTMformer
This paper focuses on multivariate time series forecasting tasks, which can be formally defined as follows: Given a historical observation matrix $\mathbf{X}\in\mathbb{R}^{L\times C}$, where $L$ denotes the input sequence length and $C$ represents the number of variables, the goal is to predict the future values $\mathbf{Y}\in\mathbb{R}^{H\times C}$, with $H$ indicating the prediction horizon.
3.1 Structure Overview
Motivated by the temporal and multivariate correlations in time series, we propose HTMformer, a general-purpose time series forecasting model. It adopts the encoder-only architecture and incorporates an HTME extractor to effectively capture temporal and multivariate features. As shown in Figure 3, the model includes an HTME extractor, a vanilla Transformer encoder layer, and a projection layer. The HTME extractor is designed to deeply extract temporal and multivariate features. After the time series is processed, the Transformer encoder layer is used to learn the relationships between multiple channels. Finally, a linear model serves as the projection layer. RevIN (Kim et al., 2022) is a normalization technique for time series forecasting, which can help models learn and generalize better. Each data batch is first normalized with RevIN before being fed into the model.
In HTMformer, showned in Figure 3 (Middle), the overall formulation for predicting the future sequence $Y\in\mathbb{R}^{H\times C}$ from the historical sequence $X\in\mathbb{R}^{L\times C}$ is as follows:
$X_{\text{in}}$ $=\text{CAT}(X,T),\qquad Y_{\text{in}}=\text{HTMEE}(X_{\text{in}}),$ (1)
$Y_{\text{out}}$ $=\text{Encode}(Y_{\text{in}}),\qquad Y=\text{Project}(Y_{\text{out}}).$
Following RevIN normalization, timestamps $T$ are appended as an additional variable to the sequence $X\in\mathbb{R}^{L\times C}$ before being fed into the model. $X_{\text{in}}\in\mathbb{R}^{L\times N}$, where $N$ denotes the multivariate dimension augmented with the timestamp. $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ denotes the hybrid temporal and multivariate embedding representations after the HTME extractor, where the embedding dimension of each variable is $D$. $Y_{\text{out}}\in\mathbb{R}^{N\times D}$ represents the predictive features produced by the Transformer encoder. Finally, a linear model serves as the projection layer to generate the final output $Y\in\mathbb{R}^{H\times C}$.
3.2 HTME Extractor
To fully capture informative representations from the raw time series, the HTME extractor incorporates two core components: a temporal feature extractor and a multivariate feature extractor, which enable the independent modeling of two dimensions. The final output is obtained by integrating the outputs of these two components. The HTME extractor takes multivariate time series as input and outputs embedding representations that fuse informative features, as illustrated in Figure 3 (Left).
Figure 3: Overall architecture of HTMformer. The HTME layer comprises two independent modules for temporal and multivariate feature extraction, directly replacing the original embeddings. It also employs inverted input to enable the attention mechanism to capture channels correlations.
3.2.1 Temporal Feature Extractor
In the temporal feature extraction module, the time series is first segmented into patches, with each patch treated as a single input token. Convolutional operations are then utilized to capture short-term temporal patterns within local receptive fields. The convolutional outputs are flattened into a high-dimensional tensor that integrates richer temporal information. Then, a linear projection layer is applied to further integrate local temporal features. The extracted features are subsequently projected along the temporal dimension to model the long-term correlations, and the tensor is reshaped to generate the final representation. Compared with conventional embedding layers, combining convolution with patch-based tokenization enables more accurate modeling of temporal contextual dependencies. By treating each channel as an independent input, interference from other features is reduced, thereby enhancing the model’s capability to capture temporal features effectively.
$N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}$ $=\operatorname{Conv}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (2)
$E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}$ $=\operatorname{Linear}\bigl(\operatorname{Flatten}\bigl(N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}\bigr)\bigr),$ (3)
$D_{\text{out}}$ $=\operatorname{Linear}\bigl(E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}\bigr).$ (4)
3.2.2 Multivariate Feature Extractor
In the multivariate feature extraction module, the time series is first partitioned into patches to accommodate potential temporal lags. The features of each patch are then flattened to aggregate local information, and a linear layer is applied to model the correlations among input variates. These learned relationships are aggregated across time through temporal convolution. The resulting output is fed into a GRU (Cho et al., 2014) network, whose gating mechanism selectively emphasizes or suppresses different inputs, thereby adaptively modulating the contribution of historical information. Finally, the features fused with multivariate information are expanded along the temporal dimension via a convolution layer to match the output of the temporal module.
$\left(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\right)$ $=\operatorname{Flatten}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (5)
$M^{N}$ $=\operatorname{GRU}\bigl(\operatorname{Linear}\bigl(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\bigr)\bigr),$ (6)
$V_{\text{out}}$ $=\operatorname{Conv}\bigl(M^{N}\bigr),$ (7)
here, $D_{\text{out}},V_{\text{out}}\in\mathbb{R}^{N\times D}$. We define a learnable fusion weight $\alpha$ to adaptively balance the contributions of the two modules to the downstream model, thereby enhancing the scalability of HTME for adaptation to datasets with diverse characteristics. The final output is the sum of the two modules:
$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$ (8)
where $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ denotes the embedded representation that fuses temporal and multivariate features. HTME is fed into a vanilla Transformer encoder to generate the predictive representations.
3.3 Encoder and Projection
The Transformer encoder layer is composed of a self-attention module and a position-wise feed-forward network. As temporal features have been fully extracted in the embedding stage, we adopt an inverted input design, which enables the self-attention mechanism to directly model global channel correlations. This inverted structure avoids forced alignment of variables within timestamps and facilitates the concatenation of temporal and multivariate dimensions. Consequently, it enables the attention mechanism to learn more appropriate sequence representations, leading to more accurate predictions. This reduces computational complexity from $\mathcal{O}(L^{2})$ in traditional methods to $\mathcal{O}(N^{2})$.
The attention mechanism projects $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ to generate query matrix $Q_{k}$, key matrix $K_{k}$, and value matrix $V_{k}$, where $d_{k}$ denotes the projection dimension for each attention head. Then, the scaled dot-product is used to derive the attention output $O_{k}\in\mathbb{R}^{N\times d_{k}}$:
$\left({O}_{k}\right)^{T}$ $=\text{Softmax}\left(\frac{Q_{k}K_{k}^{T}}{\sqrt{d_{k}}}\right)V_{k}.$ (9)
The self-attention block additionally consists of layer normalization and a position-wise feed-forward network with residual connections. It enables the model to learn cross-channel relationships and yield refined representations for each channel. iTransformer(Liu et al., 2024b) observes that, for numerical time-series forecasting tasks, a linear layer is sufficient to generate predictions, since forecasting can be realized by mapping variate-level representations without resorting to complex decoder architectures. Finally, a linear projection head serves as the decoder, mapping the learned representations to the target dimensionality and yielding the final predictions.
Unlike other models that deeply mine multidimensional correlations (Ying et al., 2021; Lai et al., 2018), our model does not explicitly model the associations between variables. Within HTME, a novel strategy is adopted to obtain high-dimensional representations that integrate multivariate information, which are subsequently mapped onto the temporal dimension. We utilize multivariate features as a complement to temporal features to learn richly informative embeddings, thereby achieving the goal of deeply extracting the hybrid temporal and multivariate characteristics of time series.
3.4 Complexity Analysis
HTMformer is composed of two primary components: the HTME extraction module and a Transformer encoder. The model’s computational complexity is primarily dictated by the Transformer encoder, as the introduction of the inverted input design results in both time and memory complexity scaling as $\mathcal{O}(N^{2})$. Compared with the state-of-the-art Transformer forecaster PatchTST (Nie et al., 2023) and the vanilla Transformer, HTMformer attains lower complexity because the number of variate tokens is fixed and independent of the lookback length, which confers clear advantages for long-term time series forecasting. Moreover, as the key component, the HTME extraction module is flexibly compatible with Transformer variants integrating efficient attention mechanisms, such as Informer or Reformer, It allows for further reductions in computational and memory costs and broadens the model’s applicability to a wider range of forecasting scenarios (see Appendix C).
4 Experiments
To evaluate both the effectiveness of the proposed HTME strategy in enhancing Transformer-based forecasters and its capacity for scalability, we apply HTME to various Transformer variants. Notably, HTME consistently and significantly improves the performance of Transformers. Moreover, we conduct a comprehensive evaluation of the proposed HTMformer on diverse time series forecasting and analysis tasks to further validate the advantages of the HTME strategy. Across various key tasks, HTMformer consistently outperforms current state-of-the-art methods. The detailed experimental configurations and implementation details can be found in Appendix D.
4.1 Experiment Settings
All models are evaluated under identical settings, following the Time-Series-Library (Wang et al., 2024b), to ensure the fairness of comparison and the reliability of the drawn conclusions.
Datasets.
We evaluated the performance of HTMformer on eight widely used datasets, including Weather, Traffic, Electricity, ETTh2 adopted in Autoformer (Wu et al., 2021), Solar-Energy proposed in LSTNet (Lai et al., 2018), and three PEMS datasets (PEMS03, PEMS04, PEMS08) investigated in SCINet (Liu et al., 2022). These datasets span diverse domains and sampling frequencies, providing a comprehensive experimental scenario for model evaluation.
Baselines and metrics.
We employ eight carefully selected state-of-the-art Transformer-based models and widely recognized non-Transformer-based models in the time series forecasting domain as baselines, encompassing iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Dlinear (Zeng et al., 2023), WPMixer (Murad et al., 2025), MultiPatchFormer (Naghashi et al., 2025), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023). Mean Squared Error (MSE) and Mean Absolute Error (MAE) are utilized as the assessment criteria.
4.2 Performance Promotion with HTME
To assess the effectiveness and scalability of HTME, it is integrated into the Transformer (Vaswani et al., 2017) and its variants, as shown in Table 1. Specifically, HTME achieves consistent improvements across different Transformer variants, with average performance gains of 35.8%, 34.3%, 43.6%, and 31.9% for Transformer, Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022), respectively. These results demonstrate that HTME captures more informative time series representations and significantly strengthens the modeling capacity of Transformer-based forecasters. More importantly, the temporal feature extraction module and the multivariate feature extraction module within the HTME layer operate as entirely decoupled components. Consequently, the multivariate representations produced by HTME can be seamlessly integrated with other embedding schemes that emphasize temporal characteristics, thereby further enhancing the applicability and versatility of HTME. The HTME framework is broadly applicable to Transformer-based forecasting architectures, allowing for flexible utilization of various attention mechanisms tailored to different application scenarios.
Table 1: Performance improvementst by HTME. Full results can be found in Appendix E.1.
Models Transformer Reformer Informer Flowformer Flashformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity Original
0.271
0.343
0.342
0.417
0.379
0.442
0.272
0.370
0.268
0.364
+HTME
0.185
0.272
0.201
0.292
0.202
0.292
0.198
0.289
0.194
0.287
Promotion
35.3%
20.6%
41.2%
29.9%
46.7%
33.9%
27.2%
19.4%
27.6%
21.1%
Weather Original
0.651
0.572
0.454
0.458
0.644
0.557
0.631
0.559
0.633
0.559
+HTME
0.254
0.278
0.257
0.276
0.254
0.280
0.254
0.279
0.253
0.280
Promotion
60.8%
51.3%
45.1%
39.9%
60.5%
49.7%
57.9%
50.0%
60.0%
49.9%
Traffic Original
0.666
0.366
0.706
0.389
0.825
0.465
0.655
0.359
0.661
0.362
+HTME
0.467
0.312
0.475
0.314
0.488
0.323
0.481
0.321
0.474
0.319
Promotion
29.8%
17.3%
32.7%
19.2%
40.8%
30.5%
26.5%
10.5%
28.2%
11.8%
Promotion Avg
35.8%
34.6%
43.6%
31.9%
33.1%
4.3 Long-term Forecasting
Long-term forecasting holds significant importance in various application domains, including meteorology, traffic management, and energy utilization. We conduct a comprehensive evaluation of the effectiveness of HTMformer in long-term forecasting. Table 2 presents the experimental forecasting results. Within each row, the lowest averaged MSE and MAE values across four prediction horizons are highlighted in red, and the second-lowest values are underscored in blue. HTMformer outperforms other models in long-term forecasting across various datasets, especially on high-dimensional datasets such as Traffic. We further provide a visualization of the long-term forecasting results produced by HTMformer (see Appendix I). HTMformer effectively captures the periodic characteristics of time series, delivers accurate trend forecasts for complex sequences, and demonstrates strong robustness against large-scale fluctuations. In conclusion, HTMformer demonstrates significant performance improvements and tangible practical benefits in long-term forecasting tasks.
Table 2: Average performance for long-term forecasting over prediction horizons $H\in\{96,192,336,720\}$ with fixed lookback $L=96$. Notably, MPFormer is short for MultiPatchFormer. Full results are listed in Appendix E.2.
Models
HTMformer
iTransformer
PatchTST
DLinear
FEDformer
MPFormer
WPMixer
TimeMixer
SegRNN
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL
0.185
0.272
0.244
0.290
0.202
0.286
0.225
0.318
0.222
0.333
0.186 0.273
0.196
0.281
0.241
0.328
0.218
0.302
Weather
0.254
0.277
0.259
0.279
0.255
0.278
0.265
0.316
0.313
0.362
0.252 0.274
0.246
0.273
0.262
0.286
0.253
0.3
Traffic
0.467 0.312
0.474
0.317
0.507
0.324
0.672
0.418
0.610
0.378
0.462
0.304
0.484
0.338
0.713
0.445
0.795
0.408
ETTh2
0.379
0.399
0.383
0.406
0.382
0.404
0.563
0.519
0.442
0.454
0.381
0.406
0.387
0.410
0.385
0.409
0.381
0.414
Solar
0.235
0.265
0.246
0.278
0.253
0.289
0.329
0.400
0.311
0.392
0.235
0.267
0.27
0.303
0.302
0.323
0.252
0.304
PEMS03
0.289
0.369
0.360
0.421
0.502
0.51
0.442
0.498
0.467
0.503
0.336 0.413
0.511
0.497
0.726
0.614
0.413
0.456
PEMS04
0.284
0.376
0.406
0.455
0.624
0.575
0.441
0.494
0.471
0.507
0.381 0.441
0.586
0.548
0.800
0.652
0.452
0.485
PEMS08
0.508
0.448
0.598
0.498
0.719
0.573
0.69
0.556
0.758
0.61
0.521 0.462
0.724
0.557
0.956
0.664
0.621
0.511
Count
6
6
0
0
0
0
0
0
0
0
3 1
1
1
0
0
0
0
4.4 Short-term Forecasting
Short-term time series forecasting tasks are also prevalent across a wide range of application domains. Table 3 presents the experimental results on the PEMS (PEMS03, PEMS04, and PEMS08) datasets. The proposed HTMformer demonstrates better performance across all three PEMS datasets. It outperforms the second-best model, MultiPatchFormer, yielding a 21.7% reduction in MSE and a 12.0% reduction in MAE, respectively. We attribute this performance gain to the exploitation of rich multivariate representations in HTME. Since short-term forecasting, constrained by the limited prediction length, does not entail intricate long temporal dependencies, it renders inter-variable correlations more salient. These results validate the robust performance of the HTMformer on short-term forecasting tasks, further complementing its strong performance in long-term forecasting.
Table 3: Short-term forecasting results, averaged across four prediction lengths $H\in\{12,24,48,96\}$ with a fixed lookback window of $L=96$. Full results are available in Appendix E.3.
Models
HTMformer
iTransformer
PatchTST
DLinear
FEDformer
MPFormer
WPMixer
TimeMixer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03
0.138
0.245
0.181
0.283
0.267
0.350
0.278
0.377
0.206
0.325
0.172 0.278
0.262
0.336
0.348
0.391
PEMS04
0.138
0.249
0.216
0.308
0.323
0.385
0.295
0.388
0.206
0.328
0.207
0.305
0.304
0.369
0.387
0.426
PEMS08
0.187
0.278
0.238
0.31
0.297
0.361
0.377
0.414
0.286
0.359
0.213 0.296
0.314
0.360
0.392
0.408
Count
3
3
0
0
0
0
0
0
0
0
0
0
0
0
0
0
Compared with iTransformer (Liu et al., 2024b), which explicitly models multivariate correlations, and PatchTST (Nie et al., 2023), which excels at capturing fine-grained temporal patterns, HTMformer achieves broadly superior performance on both long-term and short-term forecasting tasks. Even when compared with the previously best-performing Transformer-based forecaster, MultiPatchFormer, HTMformer consistently achieves either superior or comparable predictive accuracy. This performance gain is particularly pronounced on high-dimensional time series datasets, such as Traffic and Solar-Energy. Such effectiveness is attributed to the use of the HTME extractor, which facilitates multi-dimensional feature fusion in time series. HTME adaptively integrates temporal and multivariate representations to achieve comprehensive, multidimensional feature complementarity. Furthermore, it encapsulates richer and more discriminative semantic representations, enabling Transformer-based forecasters to more fully leverage sequence information. This demonstrates the broad applicability across diverse datasets, thereby highlighting the effectiveness of HTME.
5 Model Analyses
5.1 Ablation Studies
To evaluate the contribution and reliability of each module within HTMformer, we conducted ablation studies (see Appendix F). Specifically, we examined two different variants of HTMformer: 1) HTMformerV1, which utilizes only the temporal feature extraction module of the HTME extractor; 2) HTMformerV2, which utilizes only the multivariate feature extraction module of the HTME extractor. We also examined another variant 3) iTransformerV3 (iTransV3), which integrates the multivariate feature extraction module of the HTME extractor into the embedding layer of iTransformer. We adopt iTransformer (Liu et al., 2024b) as the state-of-the-art baseline. Sharing a highly similar architectural design with HTMformer, both models employ a dedicated embedding module and process inverted inputs before feeding them into the Transformer encoder. However, in iTransformer, the embedding layer conducts feature extraction exclusively along the temporal dimension. Notably, since all experiments were performed under identical configurations, the HTMformer variants are directly comparable to the other models. We also adopt MultiPatchFormer (Naghashi et al., 2025), attaining the second-best performance in various forecasting tasks, as a representative proxy for other forecasting models. Specific results are documented in Table 4.
Table 4: Results (Average) of different variants under prediction lengths $H\in\{96,192,336,720\}$. The input lookback $L$ is set to 96.
Models
HTMformer
HTMformerV1
HTMformerV2
iTransformer
iTransV3
MPFormer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL
0.185
0.272
0.184
0.273
0.275
0.357
0.244
0.290
0.194
0.288
0.186
0.273
Weather
0.254
0.277
0.253
0.278
0.262
0.290
0.259
0.279
0.252
0.267
0.252
0.274
Traffic
0.467
0.312
0.477
0.318
0.748
0.440
0.474
0.317
0.473
0.321
0.462
0.304
ETTh2
0.379
0.399
0.381
0.405
0.444
0.444
0.383
0.406
0.384
0.406
0.381
0.406
Solar
0.235
0.265
0.243
0.278
0.288
0.311
0.246
0.278
0.244
0.280
0.235
0.267
PEMS03
0.289
0.369
0.340
0.410
0.351
0.403
0.360
0.421
0.294
0.375
0.336
0.413
PEMS04
0.284
0.376
0.391
0.431
0.315
0.402
0.406
0.455
0.300
0.386
0.381
0.441
PEMS08
0.508
0.448
0.594
0.492
0.558
0.465
0.598
0.498
0.485
0.434
0.521
0.462
Effect of the temporal extractor.
Compared with other time series forecasting models, HTMformerV1 already outperforms several baselines and suffers only a marginal accuracy drop even when benchmarked against state-of-the-art approaches. Blue numbers indicate the cases where HTMformerV1 outperforms iTransformer. This observation suggests that our temporal feature extraction module has effectively captured informative temporal patterns.
Effect of the multivariate features.
To facilitate an intuitive comparison, we evaluate HTMformerV1 against HTMformerV2, with numerical values underlined when HTMformerV2 achieves better performance over HTMformerV1. Across most datasets, particularly those with a large number of variables, HTMformerV2 performs substantially worse than other time series forecasting models, highlighting that effective forecasting primarily depends on accurate temporal dependency modeling. Nevertheless, on certain datasets, HTMformerV2 attains superior or comparable results relative to most baseline models. This emphasizes that multivariate features are crucial and indispensable sources of information in time series forecasting.
Effect of the hybrid strategy.
We further conduct comparative evaluations between HTMformer and HTMformerV1, as well as between iTransformerV1 and the original iTransformer. Red numbers indicate cases where the versions augmented with multivariate integration in the embedding layer outperform their original counterparts. Across most datasets, both HTMformer and iTransformerV1 consistently achieve superior performance. This provides strong evidence for the effectiveness of our proposed strategy for incorporating multivariate features into embedding representations.
5.2 Model Efficiency
To assess the computational efficiency of HTMformer, we perform a comparative analysis of its training time, GPU memory footprint, and parameter count against state-of-the-art models, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023). All models are evaluated on the same GPU under identical lookback windows, prediction horizons, and batch sizes to ensure fair comparison. As shown in Figure 4, using the state-of-the-art MultiPatchFormer as the reference baseline, HTMformer achieves a training runtime approximately one-third of that of MultiPatchFormer, while requiring only 20% to 45% of its GPU memory footprint. Moreover, HTMformer’s model parameters are just half those of MultiPatchFormer. Importantly, these gains are achieved without sacrificing predictive accuracy, as HTMformer consistently attains superior or comparable MSE and MAE across eight benchmark datasets. These results indicate that HTMformer enables faster training and inference under the same hardware conditions, and its smaller memory footprint also enhances the model’s applicability (see Appendix G).
Figure 4: We compare HTMformer with three state-of-the-art models on eight benchmark datasets. The input sequence length is uniformly set to 96, and the prediction sequence length is 192.
It is noteworthy that the ablation studies and the long-term forecasting experiments were conducted under identical configurations, rendering the results directly comparable. We observe that for time series data, using patching and convolutions for local patterns and linear layers for long-range dependencies is generally sufficient for effective temporal modeling. Excessive modeling of temporal dependencies yields limited gains in forecasting accuracy while substantially increasing compute and memory overhead. In this study, we adopt a hybrid design in the embedding layer, employing two lightweight models to extract multidimensional features. The resulting hybrid embeddings extract richer informative representations from the sequences, thereby assisting Transformer-based forecasters in modeling time series. This approach not only improves predictive performance but also markedly reduces both computational demand and memory footprint.
6 Conclusion and Limitations
Motivated by the inherent properties of multivariate time series, we introduce the HTME strategy. The HTME module jointly encodes temporal dynamics and inter-variate dependencies present in the raw sequences, yielding embedding representations that are richer and more informative than those produced by prior designs. By inverting the input representation, the attention mechanism is empowered to more effectively model inter-channel dependencies. Experimental results demonstrate that HTMformer delivers state-of-the-art forecasting performance with faster training and a reduced GPU memory footprint, thereby validating the effectiveness of injecting multivariate information into the embedding space. Moreover, the proposed HTME module in HTMformer is seamlessly compatible with diverse Transformer architectures, underscoring its strong scalability.
Through numerous experiments, this paper further highlights that incorporating both temporal and multivariate information enables Transformer-based predictors to make more accurate predictions. The detailed discussion is provided in Appendix J. However, time series data exhibit complex correlations, and simply adding multivariate features to the temporal dimension through the embedding layer cannot fully model such a complex pattern. Jointly modeling spatiotemporal dependencies in time series remains an open problem that requires further exploration. Future improvements in Transformers can focus on devising methods to efficiently capture multivariate correlations.
7 AI Usage Statement
AI tools were used only for minor language polishing to standardize expressions. No AI assistance was employed for idea generation, data analysis, experiment design, or content creation. All intellectual contributions, results, and conclusions are the authors’ own.
8 Reproducibility Statement
All relevant implementation details, such as dataset specifications, main model hyperparameters, and experimental setups, are provided in the Appendix. The source code will be released publicly once the manuscript is accepted.
9 Ethics Statement
This study only focuses on the domain of time series forecasting. All datasets employed are widely recognized, real-world, and publicly available. Therefore, there is no potential ethical risk.
10 Acknowledgements
Appendix A Experiments of Different Embedding Methods
To demonstrate that Transformer-based forecasters have limited ability to understand sequences and the necessity of deeply mining sequence features in the embedding layer, we compared the prediction performance of Transformer (Vaswani et al., 2017), iTransformer (Liu et al., 2024b), and PatchTST (Nie et al., 2023) with different embedding approaches on four datasets: ECL, Weather, ETTh2, and Solar. The input sequence length is set to 48, 96, 144, 192, 240, and 288, with the prediction length being 96. The experimental results are shown in Figure 5.
Experimental results reveal that increasing the input sequence length initially leads to a significant decline in prediction loss. However, as the sequence length continues to grow, this benefit markedly lessens, and in some cases, prediction loss may even increase. These findings suggest that Transformer-based models exhibit limitations in proficiency in capturing effective features from raw series data. Simply increasing the lookback window does not consistently capture more informative patterns and will also incur additional computational overhead. This is because time series forecasting focuses on capturing short-term dependencies (Kim et al., 2025) and the temporal dimension inherently offers limited contextual information. Therefore, traditional Transformer-based forecasters encounter bottlenecks in reliably improving prediction accuracy. Notably, employing more sophisticated embedding methods to capture complex spatiotemporal dependencies can effectively address this problem. Motivated by these findings, we propose a novel embedding strategy that moderately extracts features along the temporal dimension to balance prediction accuracy and computational complexity, while integrating multivariate features to enhance informative content, thereby enabling the efficient acquisition of high-quality embeddings. Combined with input inversion, this approach brings comprehensive improvements to Transformer-based forecasters.
Figure 5: On the ECL, Weather, ETTh2, and Solar, we compared three models with input lengths of 48, 96, 144, 192, 240, 288 and an output length of 96. The evaluation metrics were MSE and MAE.
Appendix B Limitations of Conventional Embedding
Time series data inherently exhibits both temporal and multivariate (spatial) characteristics. A growing body of research has demonstrated the effectiveness of explicitly capturing intra- and inter-channel dependencies (Zhang & Yan, 2023; Liu et al., 2024b; a). In prevailing Transformer-based forecasting frameworks, the embedding stage typically consists of three main components: timestamp embedding, positional embedding, and time series convolution. These components are primarily designed to capture temporal patterns while often overlooking the intrinsic correlations among different variables within the data. In fact, many types of time series data exhibit significant multivariate correlations, such as traffic flow and weather forecasting (Li et al., 2018; Wu et al., 2020b), where sensors are installed at fixed spatial locations. However, conventional embeddings capture only temporal patterns, while the subsequent attention layer also primarily models temporal dependencies, thereby overlooking multivariate dependencies and leading to information loss. Due to the overemphasis on temporal modeling and the loss of multivariate information under such designs, existing Transformer-based forecasters have encountered significant bottlenecks in capturing the contextual relationships essential for accurate multivariate time series forecasting.
Furthermore, the temporal feature extraction approach of the embedding layer, inspired by large language models (LLMs), is ill-suited for time series analysis. Unlike the Transformers used in natural language and vision tasks (Kirillov et al., 2023; OpenAI, 2023), which require learning dependencies among thousands to millions of tokens, Transformers in time series forecasting generally operate within relatively limited contexts, involving no more than hundreds of time series tokens. A smaller number of tokens makes it difficult for the attention mechanism to learn clear temporal correlations. However, time series convolution is limited to extracting local patterns from the original sequence and lacks the capability for fine-grained feature extraction. Additionally, PRformer (Yu et al., 2025) highlights that traditional Transformer algorithms attempt to learn sequence variation patterns by combining positional encoding with multivariate data at different time instants. However, real-world time series are frequently non-stationary, and the assumption of translation invariance makes it difficult for the Transformer’s self-attention mechanism to learn correct temporal dependencies. Our experimental results (see Appendix A) indicate that the quality of the embeddings has a substantial impact on the performance of Transformer-based forecasters. Therefore, it is imperative to redesign the embedding model to capture more effective representations.
The HTME extractor discards positional encoding and treats timestamp information as a separate variable, concatenated to the time series inputs. This method has enhanced the model’s ability to learn the relationship between the sequence and time. It employs patch-based operations and convolution to more effectively capture short-term dependencies, and a linear layer to capture long-term correlations. This method can extract temporal features more effectively and is better suited for Transformer-based architectures in time series forecasting tasks. We further employ a well-designed multivariate feature extractor to enrich the embedding representations with multivariate features. Notably, we do not explicitly model multivariate correlations, as doing so would incur substantial computational overhead. Finally, it incorporates multivariate features into temporal features to complement and adaptively balances the contributions of the two feature components, thereby enhancing the scalability of the embeddings. The resulting HTME incorporates rich temporal information and thoroughly accounts for multivariate correlation effects. It is more effective than existing embeddings in extracting informative representations from raw time series. This approach has already achieved good results in temporal feature extraction, making it unnecessary to overemphasize temporal correlations through the attention mechanism. So HTMformer utilizes inverted input to enable the attention mechanism to capture correlations among multiple features. iTransformer (Liu et al., 2024b) demonstrates that this approach leads to improved performance.
Appendix C Time and Memory Complexity
Table 5 provides a detailed comparison of the time complexity and memory usage across various Transformer-based forecasters. $L$ denotes the input sequence length, $N$ denotes the number of variables, and $S$ denotes the stride length of PatchTST.
Table 5: Time and memory complexity of different Transformer-based forecasters.
Methods Time Memory
Transformer $O(L^{2})$ $O(L^{2})$
Informer $O(L\log L)$ $O(L\log L)$
Reformer $O(L\log L)$ $O(L\log L)$
PatchTST $O({L^{2}}/{S^{2}})$ $O({L^{2}}/{S^{2}})$
iTransformer $O(N^{2})$ $O(N^{2})$
HTMformer $O(N^{2})$ $O(N^{2})$
Appendix D Detailed Experiment Settings
D.1 Dataset Descriptions
We perform extensive experiments on eight real-world datasets to evaluate the effectiveness of the proposed HTME strategy. These datasets cover diverse domains and temporal resolutions, including:
  • ECL. Contains hourly electricity consumption data of 321 clients from 2012 to 2014.
  • Weather. Includes 21 meteorological variables collected every 10 minutes in 2020 from the meteorological station at the Max Planck Institute for Biogeochemistry.
  • Traffic. Consists of hourly road occupancy rates collected from 862 sensors on highways in the San Francisco Bay Area between January 2015 and December 2016.
  • ETTh. Comprises hourly electricity transformer temperature readings collected over two years from two counties in China, with seven recorded variables.
  • Solar-Energy. Records 10-minute solar power production records of 137 photovoltaic (PV) plants in Alabama from the year 2006.
  • PEMS. Includes PEMS03, PEMS04, and PEMS08 datasets, each containing public transportation network data collected at 5-minute intervals in California.
To ensure the consistency of our experimental approach and make sure there are no data leakage issues, we split the ETT and PEMS datasets into the training, validation, and test sets, with a 6:2:2 ratio and other datasets with a 7:1:2 ratio, following the protocols recommended in recent literature (Wu et al., 2021; Zeng et al., 2023). In all forecasting experiments, the lookback window was fixed at 96 time steps, with prediction horizons of 96, 192, 336, 720 for the long-term experiments and 12, 24, 48, 96 for the short-term experiments. The details of all datasets are provided in Table 6.
Table 6: Detailed dataset descriptions. Dim denotes the number of variables in each dataset. Dataset Size indicates the number of time points in splits respectively. Prediction Length is the number of future time points to forecast. Frequency denotes the sampling interval of the time points.
[…]
Information
ETTh2 7
{96, 192, 336, 720}
(8545, 2881, 2881) 15 min 0.45
Temperature
Electricity 321
{96, 192, 336, 720}
(18317, 2633, 5261) Hourly 0.77
Electricity
Traffic 862
{96, 192, 336, 720}
(12185, 1757, 3509) Hourly 0.68
Transportation
Weather 21
{96, 192, 336, 720}
(36792, 5271, 10540) 10 min 0.75
Weather
Solar-Energy 137
{96, 192, 336, 720}
(36601, 5161, 10417) 10 min 0.33
Solar Energy
PEMS03 358
{12, 24, 48, 96, 192, 336, 720}
(15617, 5135, 5135) 5 min 0.65
Transportation
PEMS04 307
{12, 24, 48, 96, 192, 336, 720}
(10172, 3375, 3375) 5 min 0.45
Transportation
PEMS08 170
{12, 24, 48, 96, 192, 336, 720}
(10690, 3548, 265) 5 min 0.52
Transportation
  • $\ast$ The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (Goerg, 2013). A larger value indicates better predictability.
D.2 Baseline Details
We employ five state-of-the-art Transformer-based models, such as iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), WPMixer (Murad et al., 2025), and MultiPatchFormer (Naghashi et al., 2025), as well as three well-acknowledged non-Transformer-based models, Dlinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023). We utilized the official repositories of the Time-Series-Library (Wang et al., 2024b) directly for reproduction, and adopted the same comparison settings. The Time-Series-Library consolidates and integrates implementations of prior time-series forecasting models, making it convenient for us to conduct experiments. Among these models, MultiPatchFormer is a highly advanced architecture and yields the best overall forecasting accuracy. Both iTransformer and our model utilize input inversion to enable the attention mechanism to effectively capture multivariate correlations. WPMixer efficiently captures both the frequency-domain and time-domain information in time series data, enabling deep temporal feature extraction. FEDformer achieves high forecasting accuracy and computational efficiency through an innovative architecture that integrates seasonal-trend decomposition with frequency-domain enhancement. Meanwhile, DLinear is a state-of-the-art linear model with a non-Transformer architecture. SegRNN has demonstrated outstanding performance on multiple benchmark time series forecasting datasets by employing segment-wise iterative processing and parallelized multi-step prediction. TimeMixer is a linear time series forecasting model that leverages a multi-scale decomposable mixing mechanism, demonstrating robust performance on high-dimensional datasets. Note that SegRNN is incompatible with short-term forecasting, so no comparison is conducted.
D.3 Experiment Details
All experiments were implemented in PyTorch (Paszke et al., 2019) and conducted on a dedicated NVIDIA Quadro RTX 6000 GPU. We trained the model for 10 epochs using the Adam optimizer (Kingma & Ba, 2015), with an early stopping mechanism configured at a patience value of 3. The key parameter configurations are listed in Table 7 (WPMixer is configured with d_model = 16 and d_ff = 32). We verified that, for these baselines, all other hyperparameters were taken from their official repositories within the Time-Series-Library (Wang et al., 2024b), which consolidates implementations of prior time-series forecasting models. This ensures consistency with the established fair comparison protocol. Unless otherwise stated, all models adopt the aforementioned configurations across all experiments. The only modification was to the output sequence lengths.
  • dropout: Dropout rate.
  • batch size: Batch size used for training.
  • l_rate: The learning rate used in the optimization process.
  • d_ff: Dimension of Fully FFN
  • d_model: Dimension of model
  • e_layers: Number of Transformer encoder layers.
  • n_heads: Num of heads
  • seq_len: Inputs sequence length
  • label_len: Start token length
  • loss: Loss function used for training
Table 7: Experiments of all models on all datasets follow the following experimental settings to ensure the accuracy of the experiments.
dropout batch size l_rate d_ff d_model e_layers n_heads seq_lenl label_len loss 0.1 32 0.0001 1024 512 2 8 96 48 L2
Appendix E Full Results
E.1 Performance Promotion with HTME
To assess the effectiveness and scalability of the HTME for enhancing the Transformer’s capacity to learn temporal sequence representations, we integrated HTME into the Transformer and its variants: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022). The length of historical data sequences was set to 96 time steps, with the forecast horizon varying among 96, 192, 336, and 720 time steps. The full predictive outcomes are detailed in Table 8. On three benchmark datasets (ECL, Traffic, and Weather), the HTME versions yield significant improvements over all four models across all proposed prediction horizons, thereby confirming their effectiveness and scalability. We further integrated the multivariate feature extraction module from the HTME extractor into iTransformer to assess the scalability capability of this module, with the detailed results presented in the ablation study. This component was seamlessly incorporated into the iTransformer’s embedding layer, leading to a substantial improvement in its overall performance.
Table 8: Full results of original Transformers and Transformers with HTME.
Models Transformer Reformer Informer Flowformer Flashformer
Dataset Type Step MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity Original 96 0.264 0.363 0.304 0.391 0.330 0.415 0.272 0.371 0.263 0.361
192 0.275 0.375 0.350 0.426 0.356 0.438 0.269 0.371 0.270 0.369
336 0.262 0.362 0.354 0.429 0.376 0.454 0.264 0.364 0.263 0.360
720 0.286 0.373 0.361 0.422 0.454 0.464 0.284 0.374 0.279 0.368
Avg 0.271 0.363 0.342 0.417 0.379 0.442 0.272 0.370 0.268 0.364
+HTME 96 0.157 0.247 0.169 0.266 0.170 0.266 0.170 0.266 0.163 0.262
192 0.171 0.260 0.190 0.283 0.187 0.279 0.188 0.280 0.183 0.277
336 0.192 0.277 0.200 0.293 0.207 0.296 0.202 0.294 0.199 0.292
720 0.220 0.304 0.248 0.329 0.247 0.329 0.235 0.319 0.229 0.316
Avg 0.185 0.272 0.201 0.292 0.202 0.292 0.198 0.289 0.194 0.287
Weather Original 96 0.372 0.417 0.311 0.355 0.420 0.452 0.283 0.364 0.326 0.395
192 0.632 0.580 0.453 0.464 0.423 0.455 0.548 0.541 0.587 0.544
336 0.699 0.606 0.476 0.474 0.588 0.539 0.614 0.548 0.714 0.614
720 0.903 0.688 0.579 0.541 1.146 0.784 1.079 0.783 0.906 0.684
Avg 0.651 0.572 0.454 0.458 0.644 0.557 0.631 0.559 0.633 0.559
+HTME 96 0.163 0.209 0.164 0.209 0.170 0.213 0.167 0.212 0.168 0.215
192 0.222 0.261 0.214 0.255 0.220 0.260 0.216 0.255 0.218 0.258
336 0.272 0.290 0.270 0.288 0.274 0.298 0.279 0.302 0.275 0.298
720 0.359 0.352 0.351 0.348 0.354 0.351 0.354 0.350 0.352 0.349
Avg 0.254 0.277 0.249 0.275 0.254 0.280 0.254 0.279 0.253 0.380
Traffic Original 96 0.658 0.364 0.715 0.399 0.731 0.413 0.646 0.359 0.637 0.353
192 0.655 0.359 0.705 0.390 0.761 0.431 0.646 0.355 0.659 0.364
336 0.657 0.361 0.703 0.386 0.840 0.474 0.657 0.358 0.653 0.357
720 0.697 0.380 0.703 0.383 0.969 0.545 0.671 0.364 0.697 0.375
Avg 0.666 0.366 0.706 0.389 0.825 0.465 0.655 0.359 0.661 0.362
+HTME 96 0.439 0.300 0.445 0.302 0.466 0.314 0.451 0.309 0.439 0.300
192 0.450 0.302 0.463 0.308 0.468 0.312 0.466 0.310 0.460 0.310
336 0.462 0.310 0.478 0.313 0.488 0.322 0.484 0.320 0.480 0.321
720 0.517 0.339 0.514 0.334 0.530 0.347 0.524 0.345 0.516 0.342
Avg 0.467 0.312 0.475 0.314 0.488 0.323 0.481 0.321 0.474 0.319
E.2 Long-term Forecasting Results
The results of the multivariate long-term forecasting tasks are summarized in Table 9. Lower Mean Squared Error (MSE) and Mean Absolute Error (MAE) values indicate superior predictive accuracy. The best results are highlighted in red, and the second-best results in blue. Notably, HTMformer achieves state-of-the-art performance in most evaluation scenarios, ranking first in 26 out of 40 MSE metrics and 29 out of 40 MAE metrics. Moreover, our model consistently ranks within the top two across the majority of evaluation scenarios. Table 10 also reports the standard deviations of HTMformer over five runs with different random seeds, demonstrating its performance stability.
Table 9: Full results for the long-term forecasting task. We compare extensive competitive models under different prediction lengths. Avg is averaged from all four prediction lengths.
Models
HTMformer
iTransformer
PatchTST
DLinear
FEDformer
MPFormer
WPMixer
TimeMixer
SegRNN
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Electricity
96
0.157
0.247
0.163
0.252
0.180
0.266
0.210
0.301
0.195
0.309
0.159 0.249
0.170
0.259
0.213
0.301
0.191
0.278
192
0.171
0.260
0.175 0.263
0.185
0.271
0.210
0.304
0.202
0.315
0.171
0.260
0.180
0.267
0.225
0.316
0.202
0.287
336
0.192
0.277
0.282
0.299
0.202
0.288
0.223
0.319
0.229
0.342
0.188
0.277
0.196
0.284
0.242
0.334
0.221
0.305
720
0.220
0.304
0.357
0.348
0.241
0.319
0.257
0.349
0.264
0.367
0.228 0.309
0.238
0.317
0.285
0.363
0.261
0.339
Avg
0.185
0.272
0.244
0.290
0.202
0.286
0.225
0.318
0.222
0.333
0.186 0.273
0.196
0.281
0.241
0.328
0.218
0.302
Weather
96
0.164
0.209
0.176
0.216
0.174
0.217
0.196
0.256
0.224
0.304
0.168
0.209
0.165
0.209
0.179
0.225
0.167
0.230
192
0.222
0.261
0.223
0.255
0.220
0.256
0.238
0.299
0.281
0.348
0.213
0.250
0.210
0.251
0.230
0.268
0.214
0.275
336
0.272
0.290
0.280
0.298
0.276
0.296
0.281
0.330
0.339
0.381
0.273
0.293
0.266
0.291
0.284
0.304
0.271
0.317
720
0.359
0.352
0.353 0.346 0.353 0.346
0.345
0.381
0.408
0.417
0.354
0.347
0.345
0.344
0.356
0.350
0.360
0.378
Avg
0.254
0.277
0.259
0.279
0.255
0.278
0.265
0.316
0.313
0.362
0.252 0.274
0.246
0.273
0.262
0.286
0.253
0.300
Traffic
96
0.439 0.300
0.442
0.302
0.494
0.313
0.696
0.428
0.580
0.362
0.433
0.290
0.516
0.336
0.688
0.429
0.781
0.401
192
0.450
0.302 0.459
0.308
0.490
0.330
0.646
0.407
0.606
0.379
0.450
0.296
0.512
0.331
0.687
0.434
0.789
0.405
336
0.462
0.310
0.479
0.319
0.502
0.317
0.653
0.409
0.613
0.380
0.467
0.306
0.353
0.334
0.719
0.451
0.797
0.405
720
0.517
0.339
0.516
0.342
0.545
0.337
0.694
0.428
0.641
0.393
0.500
0.326
0.558
0.353
0.761
0.466
0.813
0.421
Avg
0.467 0.312
0.474
0.317
0.507
0.324
0.672
0.418
0.610
0.378
0.462
0.304
0.484
0.338
0.713
0.445
0.795
0.408
ETTh2
96
0.300
0.348
0.295 0.344
0.293
0.343
0.341
0.395
0.350
0.391
0.299
0.350
0.293
0.345
0.296
0.345
0.295
0.355
192
0.389
0.402
0.375
0.398
0.373
0.393
0.481
0.479
0.441
0.449
0.384
0.401
0.382
0.401
0.384
0.401
0.375 0.397
336
0.409
0.408
0.436
0.435
0.422
0.430
0.592
0.542
0.498
0.490
0.411 0.428
0.424
0.437
0.429
0.439
0.414
0.436
720
0.421
0.439
0.429
0.447
0.440
0.452
0.840
0.661
0.480
0.487
0.430
0.446
0.449
0.458
0.433
0.451
0.443
0.469
Avg
0.379
0.399
0.383
0.406
0.382
0.404
0.563
0.519
0.442
0.454
0.381
0.406
0.387
0.410
0.385
0.409
0.381
0.414
Solar-Energy
96
0.199
0.239
0.213
0.253
0.214
0.257
0.289
0.377
0.279
0.363
0.203 0.240
0.232
0.278
0.256
0.292
0.232
0.294
192
0.233
0.263
0.242
0.274
0.254
0.296
0.319
0.397
0.288
0.378
0.237 0.268
0.268
0.303
0.286
0.315
0.253
0.306
336
0.252
0.279
0.262
0.290
0.284
0.314
0.352
0.415
0.316
0.399
0.249
0.280
0.291
0.316
0.323
0.337
0.264
0.310
720
0.259
0.280
0.270
0.296
0.263
0.291
0.356
0.412
0.363
0.430
0.254
0.283
0.290
0.315
0.346
0.349
0.260
0.306
Avg
0.235
0.265
0.246
0.278
0.253
0.289
0.329
0.400
0.311
0.392
0.235
0.267
0.270
0.303
0.302
0.323
0.252
0.304
PEMS03
96
0.25
0.343
0.337
0.413
0.504
0.516
0.458
0.517
0.325
0.423
0.302 0.395
0.514
0.505
0.731
0.625
0.381
0.438
192
0.297
0.378
0.374
0.436
0.547
0.544
0.477
0.527
0.475
0.515
0.349 0.428
0.557
0.526
0.797
0.648
0.447
0.478
336
0.274
0.356
0.332
0.395
0.432
0.460
0.396
0.456
0.425
0.480
0.313 0.388
0.440
0.450
0.614
0.545
0.380
0.432
720
0.336
0.401
0.397
0.440
0.527
0.521
0.439
0.492
0.644
0.596
0.383
0.441
0.533
0.509
0.762
0.638
0.447
0.478
Avg
0.289
0.369
0.360
0.421
0.502
0.510
0.442
0.498
0.467
0.503
0.336 0.413
0.511
0.497
0.726
0.614
0.413
0.456
PEMS04
96
0.226
0.335
0.387
0.446
0.645
0.594
0.452
0.504
0.300 0.411
0.356
0.428
0.584
0.554
0.794
0.658
0.415
0.468
192
0.286
0.383
0.429
0.473
0.688
0.615
0.477
0.527
0.458
0.513
0.398 0.459
0.650
0.586
0.891
0.697
0.496
0.510
336
0.286
0.372
0.369
0.426
0.536
0.516
0.396
0.456
0.411
0.474
0.348 0.413
0.511
0.499
0.703
0.596
0.420
0.461
720
0.338
0.414
0.440
0.476
0.629
0.578
0.439
0.492
0.718
0.630
0.423 0.467
0.601
0.556
0.812
0.660
0.479
0.502
Avg
0.284
0.376
0.406
0.455
0.624
0.575
0.441
0.494
0.471
0.507
0.381 0.441
0.586
0.548
0.800
0.652
0.452
0.485
PEMS08
96
0.365
0.409
0.477
0.470
0.567
0.534
0.672
0.564
0.504
0.506
0.406 0.437
0.648
0.556
0.854
0.657
0.516
0.487
192
0.513
0.462
0.627
0.518
0.794
0.620
0.726
0.580
0.726
0.608
0.539 0.479
0.766
0.581
1.048
0.702
0.646
0.528
336
0.530
0.435
0.599
0.472
0.709
0.539
0.650
0.518
0.803
0.615
0.538 0.444
0.696
0.518
0.906
0.615
0.617
0.490
720
0.624
0.487
0.691
0.532
0.807
0.599
0.713
0.562
1.000
0.711
0.603
0.489
0.789
0.575
1.019
0.682
0.707
0.541
Avg
0.508
0.448
0.598
0.498
0.719
0.573
0.690
0.556
0.758
0.610
0.521 0.462
0.724
0.557
0.956
0.664
0.621
0.511
Count
26
29
0
0
2
2
0
0
0
0
10 9
6
3
0
0
0
0
Table 10: Robustness of HTMformer performance is evaluated over five random seeds.
Dataset ECL Weather Traffic ETTh2 Horizon MSE MAE MSE MAE MSE MAE MSE MAE 96 $0.157\pm 0.003$ $0.247\pm 0.002$ $0.164\pm 0.004$ $0.209\pm 0.002$ $0.439\pm 0.002$ $\pm 0.002 0.300±0.003$$\pm 0.003$ $0.348\pm 0.004$ 192 $0.171\pm 0.003$ $0.260\pm 0.002$ $0.222\pm 0.005$ $0.261\pm 0.004$ $0.450\pm 0.002$ $0.302\pm 0.003$ $0.389\pm 0.004$ $0.402\pm 0.003$ 336 $0.192\pm 0.005$ $0.277\pm 0.005$ $0.272\pm 0.005$ $0.290\pm 0.010$ $0.462\pm 0.004$ $0.310\pm 0.005$ $0.409\pm 0.004$ $0.408\pm 0.004$ 720 $0.220\pm 0.006$ $0.304\pm 0.007$ $0.359\pm 0.004$ $0.352\pm 0.002$ $0.517\pm 0.009$ $0.339\pm 0.007$ $0.421\pm 0.007$ $0.439\pm 0.007$
Dataset Solar-Energy PEMS03 PEMS04 PEMS08 Horizon MSE MAE MSE MAE MSE MAE MSE MAE 96 $0.199\pm 0.004$ $0.239\pm 0.003$ $0.250\pm 0.003$ $0.343\pm 0.002$ $0.226\pm 0.008$ $0.335\pm 0.004$ $0.365\pm 0.003$ $0.409\pm 0.005$ 192 $0.233\pm 0.004$ $0.263\pm 0.006$ $0.297\pm 0.005$ $0.378\pm 0.010$ $0.286\pm 0.010$ $0.383\pm 0.005$ $0.513\pm 0.005$ $0.462\pm 0.006$ 336 $0.252\pm 0.006$ $0.279\pm 0.007$ $0.274\pm 0.010$ $0.356\pm 0.009$ $0.286\pm 0.010$ $0.372\pm 0.005$ $0.530\pm 0.006$ $0.435\pm 0.010$ 720 $0.259\pm 0.006$ $0.280\pm 0.009$ $0.336\pm 0.007$ $0.401\pm 0.008$ $0.338\pm 0.009$ $0.414\pm 0.007$ $0.624\pm 0.003$ $0.487\pm 0.006$
E.3 Short-term Forecasting Results
Table 11 presents the short-term forecasting results. The evaluation metrics comprise the Mean Squared Error (MSE) and Mean Absolute Error (MAE) computed over four prediction horizons. Our approach consistently achieves substantial improvements across all horizons on the selected datasets. In particular, on the PEMS04 dataset, our proposed model achieved a 33.3% reduction in average MSE and an 18.3% reduction in average MAE compared to the second-best results. The standard deviations, computed over five independent runs with different random seeds, are reported in Table 12. For short-term forecasting, the limited length of the input sequence constrains the model’s capacity to capture sufficient temporal dependencies. Consequently, modeling multivariate correlations becomes crucial, in which HTMformer demonstrates superior performance.
Table 11: Full results for the short-term forecasting task. We compare extensive competitive models on PEMS datasets. Avg means the average results from all four prediction lengths.
Models
HTMformer
iTransformer
PatchTST
DLinear
FEDformer
MPFormer
WPMixer
TimeMixer
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
PEMS03
12
0.067
0.174
0.075 0.184
0.102
0.217
0.122
0.245
0.125
0.251
0.089
0.199
0.093
0.204
0.107
0.220
24
0.093
0.204
0.115
0.229
0.187
0.298
0.201
0.320
0.152
0.279
0.114 0.225
0.151
0.263
0.185
0.292
48
0.151
0.261
0.195
0.305
0.278
0.369
0.334
0.428
0.222
0.347
0.186 0.296
0.292
0.373
0.375
0.428
96
0.241
0.341
0.340
0.417
0.504
0.516
0.458
0.517
0.325
0.423
0.300 0.393
0.514
0.505
0.731
0.625
Avg
0.138
0.245
0.181
0.283
0.267
0.350
0.278
0.377
0.206
0.325
0.172 0.278
0.262
0.336
0.348
0.391
PEMS04
12
0.076
0.183
0.095 0.202
0.112
0.231
0.147
0.272
0.136
0.263
0.109
0.218
0.111
0.223
0.126
0.239
24
0.097
0.209
0.140
0.249
0.187
0.301
0.224
0.340
0.156
0.284
0.138 0.248
0.181
0.289
0.208
0.312
48
0.146
0.263
0.238
0.333
0.355
0.422
0.356
0.437
0.226
0.351
0.226 0.326
0.343
0.411
0.422
0.456
96
0.233
0.342
0.394
0.449
0.638
0.587
0.453
0.504
0.308
0.417
0.358 0.430
0.584
0.554
0.794
0.658
Avg
0.138
0.249
0.216
0.308
0.323
0.385
0.295
0.388
0.206
0.328
0.207 0.305
0.304
0.369
0.387
0.426
PEMS07
12
0.081
0.187
0.086 0.190
0.108
0.225
0.152
0.274
0.175
0.273
0.099
0.205
0.103
0.212
0.117
0.230
24
0.120
0.227
0.135
0.240
0.173
0.284
0.246
0.351
0.212
0.307
0.127 0.234
0.174
0.279
0.196
0.302
48
0.198
0.297
0.247
0.333
0.341
0.404
0.438
0.469
0.296
0.375
0.219 0.312
0.332
0.396
0.402
0.443
96
0.352
0.401
0.486
0.478
0.567
0.534
0.672
0.564
0.463
0.481
0.410 0.434
0.648
0.556
0.854
0.657
Avg
0.187
0.278
0.238
0.310
0.297
0.361
0.377
0.414
0.286
0.359
0.213 0.296
0.314
0.360
0.392
0.408
Count
15
15
0
0
0
0
0
0
0
0
0
0
0
0
0
0
Table 12: Results on short-term time series forecasting are obtained from five random seeds.
Dataset PEMS03 PEMS04 PEMS08 Horizon MSE MAE MSE MAE MSE MAE 12 $0.067\pm 0.000$ $174\pm 0.000 0.0$$076\pm 0.000$ $183\pm 0.000 0.0$$081\pm 0.002$ $0.187\pm 0.002$ 24 $0.093\pm 0.001$ $204\pm 0.000 0.0$$097\pm 0.001$ $0.209\pm 0.002$ $0.120\pm 0.001$ $0.227\pm 0.001$ 48 $0.151\pm 0.003$ $0.261\pm 0.001$ $0.146\pm 0.004$ $0.263\pm 0.004$ $0.198\pm 0.002$ $0.297\pm 0.002$ 96 $0.241\pm 0.000$ $0.341\pm 0.003$ $0.233\pm 0.005$ $0.342\pm 0.005$ $0.352\pm 0.006$ $0.401\pm 0.002$
Appendix F Additional Ablation Studies
To investigate the functional rationale of HTMformer components, we conduct detailed ablation experiments involving both component replacement and component removal. The detailed results are reported in Table 13 of the main text. As all experiments were conducted under identical configurations, the HTMformer variants are directly comparable to other models.
Effect of the temporal feature extraction module.
By comparing iTransformer and HTMformerV1, blue numbers are used to denote cases where HTMformerV1 outperforms iTransformer. Observations indicate that HTMformerV1 demonstrates improvements in 63/80 cases. By comparing HTMformerV1 and HTMformerV2, numbers prefixed with an underline indicate that HTMformerV1 outperforms HTMformerV2. Notably, HTMformerV1 performs better in 57/80 cases. Significantly, on Solar and PEMS datasets lacking timestamps, HTMformerV2 often exhibits superior performance, verifying our first hypothesis: 1) The majority of time series features are stored in the temporal dimension, and the depth of temporal feature mining substantially influences prediction accuracy.
Effect of the multivariate features.
Using iTransformer and HTMformerV1 as baselines, we compare them against iTransformerV3 and HTMformer versions. Red highlighted numbers indicate cases where the multivariate-integrated versions outperform their original counterparts. Experimental results show that iTransformerV3 achieves superior performance in 61/80 cases compared to iTransformer, while HTMformer outperforms its baseline in 68/80 cases. The gains are particularly evident on the PEMS datasets, which are derived from public transportation network data characterized by strong correlations among sensor nodes. This finding corroborates our second hypothesis: 2) The spatial dimension (multivariate dimension) also contains abundant features, and comprehensive multivariate feature mining can further reduce errors.
Effect of the hybrid strategy.
Among all models and their variants, HTMformer excels in 68/80 cases. This demonstrates the effectiveness of the combination. In most cases, temporal features are the primary characteristics of time series data, but the correlations between multiple variables cannot be overlooked. It is of great significance for improving the accuracy of Transformer-based predictors. Different datasets vary in the depth of temporal and multivariate features. Therefore, we use a learnable parameter to control the proportion of the two modules, allowing the fusion of multivariate features to still reduce errors to an extent. In conclusion, the strategy of fusing temporal features and multivariate features is superior to using either of them alone. This strategy can enhance the accuracy and scalability of Transformer-based predictors.
Table 13: We compare different variants under different prediction lengths on multiple datasets (MSE and MAE). The input sequence length is set to 96.
Models
HTMformer
HTMformerV1
HTMformerV2
iTransformer
iTransV3
MPFormer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 96 0.157 0.247 0.153 0.246 0.255 0.34 0.163 0.252 0.165 0.262 0.159 0.249
192 0.171 0.260 0.169 0.260 0.264 0.349 0.175 0.263 0.184 0.278 0.171 0.260
336 0.192 0.277 0.187 0.277 0.273 0.357 0.282 0.299 0.199 0.294 0.188 0.277
720 0.220 0.304 0.227 0.311 0.31 0.384 0.357 0.348 0.231 0.318 0.228 0.309
AVG 0.185 0.272 0.184 0.273 0.275 0.357 0.244 0.29 0.194 0.288 0.186 0.273
Weather 96 0.164 0.209 0.169 0.212 0.175 0.229 0.176 0.216 0.167 0.167 0.168 0.209
192 0.222 0.261 0.218 0.256 0.223 0.264 0.223 0.255 0.217 0.256 0.213 0.250
336 0.272 0.290 0.275 0.298 0.287 0.31 0.28 0.298 0.272 0.297 0.273 0.293
720 0.359 0.352 0.353 0.349 0.364 0.358 0.357 0.349 0.353 0.349 0.354 0.347
AVG 0.254 0.277 0.253 0.278 0.262 0.29 0.259 0.279 0.252 0.267 0.252 0.274
Traffic 96 0.439 0.300 0.447 0.303 0.648 0.405 0.442 0.302 0.440 0.309 0.433 0.290
192 0.450 0.302 0.460 0.307 0.737 0.438 0.459 0.308 0.457 0.305 0.450 0.296
336 0.462 0.310 0.482 0.319 0.745 0.435 0.479 0.319 0.478 0.321 0.467 0.306
720 0.517 0.339 0.522 0.343 0.863 0.484 0.516 0.342 0.520 0.349 0.500 0.326
AVG 0.467 0.312 0.477 0.318 0.748 0.44 0.474 0.317 0.473 0.321 0.462 0.304
ETTh2 96 0.300 0.348 0.298 0.348 0.375 0.399 0.295 0.344 0.304 0.352 0.299 0.350
192 0.389 0.402 0.387 0.402 0.45 0.435 0.375 0.398 0.391 0.406 0.384 0.401
336 0.409 0.408 0.416 0.427 0.481 0.469 0.436 0.435 0.419 0.428 0.411 0.428
720 0.421 0.439 0.426 0.443 0.471 0.474 0.429 0.447 0.422 0.438 0.430 0.446
AVG 0.379 0.399 0.381 0.405 0.444 0.444 0.383 0.406 0.384 0.406 0.381 0.406
Solar 96 0.199 0.239 0.209 0.252 0.244 0.284 0.213 0.253 0.207 0.249 0.203 0.240
192 0.233 0.263 0.238 0.274 0.277 0.306 0.242 0.274 0.240 0.279 0.237 0.268
336 0.252 0.279 0.261 0.292 0.323 0.326 0.262 0.29 0.256 0.290 0.249 0.280
720 0.259 0.280 0.265 0.295 0.311 0.329 0.27 0.296 0.275 0.303 0.254 0.283
AVG 0.235 0.265 0.243 0.278 0.288 0.311 0.246 0.278 0.244 0.280 0.235 0.267
PMES03 96 0.250 0.343 0.304 0.393 0.302 0.363 0.337 0.413 0.240 0.340 0.302 0.395
192 0.297 0.378 0.353 0.424 0.373 0.422 0.374 0.436 0.300 0.384 0.349 0.428
336 0.274 0.356 0.318 0.388 0.32 0.388 0.332 0.395 0.286 0.365 0.313 0.388
720 0.336 0.401 0.387 0.438 0.409 0.44 0.397 0.44 0.352 0.414 0.383 0.441
AVG 0.289 0.369 0.340 0.410 0.351 0.403 0.36 0.421 0.294 0.375 0.336 0.413
PMES04 96 0.226 0.335 0.364 0.433 0.236 0.348 0.387 0.446 0.235 0.344 0.356 0.428
192 0.286 0.383 0.406 0.406 0.326 0.413 0.429 0.473 0.307 0.395 0.398 0.459
336 0.286 0.372 0.359 0.417 0.312 0.398 0.369 0.426 0.297 0.378 0.348 0.413
720 0.338 0.414 0.436 0.471 0.387 0.451 0.44 0.476 0.362 0.428 0.423 0.467
AVG 0.284 0.376 0.391 0.431 0.315 0.402 0.406 0.455 0.300 0.386 0.381 0.441
PMES08 96 0.365 0.409 0.469 0.461 0.401 0.416 0.477 0.47 0.338 0.388 0.406 0.437
192 0.513 0.462 0.612 0.507 0.567 0.481 0.627 0.518 0.487 0.444 0.539 0.479
336 0.530 0.435 0.606 0.471 0.567 0.452 0.599 0.472 0.516 0.425 0.538 0.444
720 0.624 0.487 0.692 0.531 0.697 0.512 0.691 0.532 0.600 0.479 0.603 0.489
AVG 0.508 0.448 0.594 0.492 0.558 0.465 0.598 0.498 0.485 0.434 0.521 0.462
Appendix G Details of Model Efficiency
For a comprehensive efficiency comparison, we evaluate HTMformer against three highly competitive baselines, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023), across eight datasets. The evaluation considers three key metrics: training time, GPU memory footprint, and total parameter count, with the input sequence length fixed at 96 and the prediction length at 192. Moreover, the experimental hardware conditions and parameter configurations follow those described in Appendix B, and the best results are marked in red. The full results are shown in Table 14 (Left) and Table 14 (Right). The training time and GPU usage serve as an intuitive indicator of total parameter count, so detailed reporting is omitted. Across all datasets, HTMformer achieves the best training time and GPU memory footprint. Notably, for high-dimensional datasets, particularly when the number of variate dimensions exceeds one hundred, HTMformer delivers nearly three times the training speed of the second-best model and requires only one-third of its memory footprint. Moreover, this advantage becomes increasingly pronounced as dimensionality grows. This superiority stems from our lightweight design philosophy, in which HTME avoids overemphasizing temporal feature extraction and HTMformer adopts only a vanilla Transformer encoder combined with the inverted input strategy.
[…]
Datasets HTMFormer MPFormer WPMixer PatchTST
Models Ours 2025 2025 2023
ECL 0.1004 0.3319 0.4143 0.5088
Weather 0.0282 0.0320 0.0392 0.0292
Traffic 0.2682 0.9648 1.2374 1.0294
ETTh2 0.0199 0.0279 0.0183 0.0190
Solar 0.0502 0.1435 0.1745 0.2200
pems03 0.0982 0.3868 8344.82 0.5797
pems04 0.1514 0.5801 0.3894 0.4857
pems07 0.0915 0.3108 0.2154 0.2552
AVG 0.1010 0.3472 0.3849 0.3909
Datasets HTMFormer MPFormer WPMixer PatchTST
Models Ours 2025 2025 2023
ECL 1623.47 7471.94 13784.07 7195.99
Weather 151.83 584.33 651.45 538.76
Traffic 6221.31 20860.23 29560.07 19152.58
ETTh2 102.48 265.89 455.10 430.54
Solar 647.04 3219.13 5970.26 3130.73
pems03 1849.57 8344.82 8015.67 8015.67
pems04 1536.60 7141.78 13195.91 6882.86
pems07 778.26 3979.58 7369.67 3869.06
AVG 1613.82 6483.46 10125.27 6152.02
Appendix H Hyperparameter Sensitivity
We assess the sensitivity of HTMformer to variations in key hyperparameters, including the learning rate $lr$, the number of Transformer blocks $L$, the FCN dimension $K$, and the hidden dimension $D$ of the variate tokens. This study can provide guidance for hyperparameter selection of HTMformer in practical applications. As shown in Figure 6, our analysis leads to the following key observations: Increasing hyperparameter values does not necessarily translate into improved forecasting accuracy. The model attains its best performance on most datasets when configured with $lr=0.0005$, $L=2$, $K=1024$, and $D=512$.
Figure 6: Hyperparameter sensitivity with respect to lrlr, $L$, $K$, and $D$ . All results are obtained using a lookback window of $T=96$ time steps and a forecast horizon of $S=96$ time steps.
Appendix I Visualization of Prediction Results
We visualize HTMformer’s ability to predict trends across various time series datasets, including ECL, Traffic, Solar-Energy, Weather, and PEMS (PEMS03, PEMS04 and PEMS08), as shown in Figures˜7, 8, 9, 10, 11, 12, 13 and 14. Each example uses a 96-step input to generate 96-step predictions. In the visualizations, the orange lines indicate the ground truth values, and the blue lines show the model’s predictions. It accurately captures the cyclical patterns and oscillatory behaviors, and successfully forecasts the overall directional trends.
To assess the performance of various models, we perform a qualitative comparison by visualizing the final forecasting results derived from two representative datasets (Weather and ETTh2). Among the various models, HTMformer exhibits superior or comparable performance in predicting the most series variations. Prediction showcases are listed in Figure 15 and Figure 16.
Appendix J Discussion on Transformer-based Predictors
Recent studies (Lu et al., 2023; Li et al., 2023b) have found that Channel Dependence ideally gains from higher capacity, while Channel Independence can significantly enhance performance due to sample scarcity. In time series forecasting, an increasing number of models employ Channel Independence (Nie et al., 2023), as most current forecasting benchmarks are not sufficiently large. These models treat time series variates independently and utilize a shared backbone, achieving good performance. However, Channel Independence inevitably overlooks the multivariate correlations that exist in most datasets, resulting in the loss of some information. We believe that making variates independent is desirable, but multivariate correlations should not be ignored. However, previous improvements to models have mainly focused on extracting features in the temporal dimension. Excessive extraction of temporal features does not further improve prediction accuracy, but rather significantly increases computational overhead.
Notably, as demonstrated in both the introduction and ablation studies of this paper. The main reason why traditional Transformer-based predictors encounter performance bottlenecks is that the embedding model extracts only limited effective information from the sequence, which in turn affects the performance of subsequent models. Due to the inherent limitations of Transformer-based predictors, simply enlarging the input sequence length cannot further enhance accuracy. Constructing an embedding model that deeply extracts sequence information can effectively address this problem. Therefore, we propose a novel strategy, specifically extracting multivariate features in the embedding layer and using them to complement temporal features. This approach incorporates multivariate correlations, thereby further extracting effective information from the sequence. Moreover, it does not affect the subsequent attention mechanism to utilize Channel Independence.
We find that extracting local patterns in time series through patching and convolution, together with capturing long-term dependencies using linear layers, is sufficient for effectively modeling temporal features. For Transformer-based predictors, this is sufficient to extract temporal representations, and it is unnecessary to employ attention mechanisms to model temporal correlations.
Graph Convolutional Networks (GCNs) typically construct an adjacency matrix to represent multivariate correlations (Kipf & Welling, 2017; Bruna et al., 2014). However, the adjacency matrix is often non-learnable. Although the physical locations of sensors are relatively fixed, the correlations among variables are not necessarily constant, which severely affects the performance of GCNs. Introducing a learnable adjacency matrix (Wang et al., 2019) can substantially increase computational overhead and lead to model instability (Wu et al., 2020a; Franceschi et al., 2019). Therefore, this paper adopts a "weak learning" strategy, whereby multivariate correlations are simply extracted and incorporated as features into tokens to correct temporal features.
We validate the effectiveness and scalability of the proposed HTME strategy by integrating it into various Transformer variants. HTME can be seamlessly combined with various attention mechanisms across a wide spectrum of time series forecasting tasks to achieve optimal performance. Our model demonstrates superior performance in short-term sequence forecasting. On the selected datasets, PEMS03, PEMS04, and PEMS08, our model achieves the best results compared with the second-best methods. In long-term sequence forecasting, our model ranks first or second on eight selected datasets in most cases. This demonstrates the effectiveness of the HTME strategy in advancing time series forecasting. We also compared the Inference Time, GPU Memory Footprint, and Model Parameters of HTMformer with those of three models that have good forecasting performance: MultiPatchFormer, WPMixer, and PatchTST. On the selected datasets, HTMformer achieves superior or equivalent performance with only about 0.2-0.45 times the memory of the three models, and is two to three times faster. These results suggest that the HTME model is highly flexible for deployment across a broader spectrum of devices and can better accommodate real-time forecasting tasks, thereby creating greater potential for practical applications and future scalability.
Figure 7: Examples of forecasts for the ECL dataset with a 96-step predictions.
Figure 8: Examples of forecasts for the Traffic dataset with a 96-step predictions.
Figure 9: Examples of forecasts for the Weather dataset with a 96-step predictions.
Figure 10: Examples of forecasts for the ETTh2 dataset with a 96-step predictions.
Figure 11: Examples of forecasts for the Solar dataset with a 96-step predictions.
Figure 12: Examples of forecasts for the PEMS03 dataset with a 96-step predictions.
Figure 13: Examples of forecasts for the PEMS04 dataset with a 96-step predictions.
Figure 14: Examples of forecasts for the PEMS08 dataset with a 96-step predictions.
Figure 15: Visualization of results on the Weather dataset across all selected models.
Figure 16: Visualization of results on the ETTh2 dataset across all selected models.

### v3

@@
 Abstract
-Transformer-based methods have achieved impressive results in time series forecasting.
-However, existing Transformers still exhibit limitations in sequence modeling as they tend to overemphasize temporal dependencies.
-This incurs additional computational overhead without yielding corresponding performance gains.
-We find that the performance of Transformers is highly dependent on the embedding method used to learn effective representations.
-To address this issue, we extract multivariate features to augment the effective information captured in the embedding layer, yielding multidimensional embeddings that convey richer and more meaningful sequence representations.
-These representations enable Transformer-based forecasters to better understand the series.
-Specifically, we introduce Hybrid Temporal and Multivariate Embeddings (HTME).
-The HTME extractor integrates a lightweight temporal feature extraction module with a carefully designed multivariate feature extraction module to provide complementary features, thereby achieving a balance between model complexity and performance.
-By combining HTME with the Transformer architecture, we present HTMformer, leveraging the enhanced feature extraction capability of the HTME extractor to build a lightweight forecaster.
-Experiments conducted on eight real-world datasets demonstrate that our approach outperforms existing baselines in both accuracy and efficiency.
+Existing predictors primarily focus on modeling the temporal dimension, often achieving superior performance even with channel-independent designs.
+This is because, due to the inherent limitations of time series data, existing multivariate feature extraction strategies tend to introduce substantial noise and brings additional computational overhead.
+However, these predictors neglect the inherent multivariate correlations within time series, inevitably leading to an inability to accurately identify complex patterns.
+To address this issue, we propose a novel multivariate extraction strategy, Hybrid Temporal and Multivariate Embedding (HTME).
+HTME separates the multivariate feature space for denoising and pattern extraction, and maps the features to the temporal feature space, yielding multidimensional embeddings that convey richer and more meaningful sequence representations.
+This strategy enables predictors focusing on the temporal dimension fully consider multivariate features with minimal additional computational cost, consistently improving their performance.
+We construct HTME-versions for eight predictors across three different predictive architectures.
+In all experiments, the resulting HTMPredictors consistently achieve state-of-the-art performance.
 1 Introduction
 Figure 1: Performance (MSE) of HTMformer. Smaller image area yields better performance.
-Long-term time series forecasting holds significant importance in various fields such as finance and economics (Andersen et al., 2006), climate science (Mudelsee, 2019), healthcare (Zeger et al., 2006), geophysics (Gubbins, 2004), industrial monitoring (Truong et al., 2022), etc.
-Recently, Transformer (Vaswani et al., 2017) and its variants have achieved tremendous success in time series forecasting, showing great promise.
-Current Transformer-based forecasters often overemphasize modeling temporal dependencies.
-However, the effective information within the temporal dimension is limited.
-Consequently, existing models incur increasing computational overhead without significant accuracy improvements.
-Unlike natural language sequences, time series data, particularly in domains such as transportation, meteorology, and electricity systems, are influenced by specific physics-based systems, e.g., traffic flow is largely governed by the real-world road network.
-Therefore, multivariate correlations are essential for time series forecasting tasks (Zheng & Sun, 2024).
-Two key points are worth noting: 1) the vast majority of features in time series are contained within the temporal dimension, and the depth of exploration in this dimension significantly influences prediction accuracy (Hyndman & Athanasopoulos, 2021); 2) the multivariate dimension also contains a large number of features (Pravilovic et al., 2013), and fully exploiting multivariate features can further boost performance.
-The performance of Transformer-based forecasters is highly dependent on the design of the embedding layer.
-Before applying multi-head self-attention to the embeddings, effective feature extraction is critical to enhancing the model’s sequence modeling capacity (Li et al., 2023a).
-We compare the ground-truth values and the predictions obtained under different embedding strategies on four datasets (see Appendix A).
-Specifically, models such as iTransformer (Liu et al., 2024b) and PatchTST (Nie et al., 2023), which employ an inverted embedding strategy or patch-based embedding strategy for deeper time-series representation learning, achieve significantly better performance than the vanilla Transformer, which only uses simple convolutional embeddings.
-Meanwhile, we also notice that as the input length increases, the prediction loss fails to markedly decrease and even tends to rise.
-These suggest that the modeling capacity of Transformer-based forecasters is inherently constrained by the quality of embeddings, and this limitation cannot be eliminated by increasing the lookback length.
-To mitigate this issue, this paper proposes to jointly capture temporal and multivariate features in the embedding layer, thereby yielding semantically rich embedding representations to overcome the inherent limitations of the conventional embedding strategy (see Appendix B).
-Utilizing proposed strategy, we constructed a Transformer-based forecaster and achieved state-of-the-art empirical performance on benchmarks, as shown in Figure 1.
+Long-term time series forecasting holds significant importance in various fields such as finance and economics (Andersen2006), climate science (Mudelsee2019), healthcare (Zeger2006), geophysics (Gubbins2004), industrial monitoring (Truong2022), etc.
+The key to time series forecasting lies in capturing temporal dependencies and multivariate correlations, while also meeting the demands of real-time prediction (Lan2022IntraInter; Yu2024Revitalizing).
+Due to the inherent constraints of time series, the extraction of multivariate features inevitably introduces significant noise(liu2024itransformer; han2023capacity; li2023revisiting).
+So existing prediction models, such as PatchTST and DLinear, focusing solely on the temporal dimension have superior performance.
+We point that this strategy of neglecting multivariate correlations inevitably prevents the model from fully capturing complex time-series patterns.
+The absence of key patterns inevitably prevents the model from effectively fitting diverse and complex time series, which severely limits its accuracy and robustness.
+Current hybrid stragetys, such as Crossformer, employing a flawed approach to extracting multivariate features.
+These models introduce extra, complex modules into their backbone networks, which significantly increases computational overhead.
+Furthermore, severe noise interference due to a shared feature space, leading to a degradation in accuracy (see Appendix A).
+Therefore, it is necessary to redesign a strategy for extracting multivariate correlations, enabling these models to reconsider them.
+To mitigate this issue, this paper proposes a strategy that hybrid temporal and multivariate features in the embedding (HTME) layer , thereby yielding semantically rich embedding representations to overcome the inherent limitations (see Appendix B).
+HTME employs parallel and specialized feature extraction modules to separately capture temporal and multivariate features from the raw input, thereby minimizing interference between them.
+More importantly, the two feature spaces in HTME are disjoint, which confines the noise exclusively to the multivariate feature space.
+Then, HTME projects the multivariate features into the temporal feature space.
+HTME is designed to leverage multivariate features to appropriately supplement the temporal feature space, yielding semantically richer embedding representations.
+This enables subsequent modules to sufficiently extract both temporal patterns and mutilvariate correlations by focusing on just a single dimension without additional moudels.
+Rather than mixing or concatenating the feature spaces, HTME suppresses noise within the multivariate space while preserves the integrity and consistency of the temporal feature space.
+HTME consistently enhances the performance of various predictors focusing on temporal dimension with only negligible computational overhead.
+Leveraging this design, we construct a novel Transformer-based framework(HTMformers), as illustrated in Figure 2.
+HTMformers consistently enhance the modeling capabilities of various attention mechanisms.
+By applying the HTME strategy, even simple models can achieve performance comparable to complex models.
+HTMformer using Full-Attention as a presentative of HTMformers achieves state-of-the-art performance on real-world benchmarks, as shown in Figure 1.
+We further applied the HTME strategy to other forecasting models with different architectures, and it consistently improved their performance.This demonstrates the validity and strong generalizability of the HTME strategy.
 The contributions of this work are as follows:
-  • We propose the Hybrid Temporal and Multivariate Embedding (HTME) extractor, which comprises a lightweight temporal feature extractor to capture informative temporal dynamics and a multivariate feature extractor to model the multivariate dependencies.
-The HTME is broadly applicable across Transformer-based forecasting architectures and effectively enhances the capacity of transformer-based models to model complex time series.
-  • We introduce a novel forecasting model, HTMformer, to further validate the effectiveness of the HTME strategy.
-In this framework, the HTME serves as input to a vanilla encoder-only Transformer framework.
-Extensive empirical evaluations are conducted on eight real-world benchmarks spanning application domains.
-Notably, HTMformer consistently achieves state-of-the-art performance across various benchmarks.
+  • We propose the Hybrid Temporal and Multivariate Embedding (HTME) strategy.
+It enables predictors focusing on the temporal dimension to fully account for multivariate features, while introducing only minimal computational overhead.
+  • We introduce a novel forecasting framework, HTMformers, which consistently improves the performance of attention mechanisms in forecasting.
+The representation HTMformer consistently achieves state-of-the-art performance across various benchmarks.
+  • We selected seven widely-used forecasting models and developed their HTME-augmented versions.
+Extensive experiments demonstrate that HTME consistently enhances the ability of diverse architectures to model complex time series, particularly for Transformer framework.
+Figure 2: Previous works have focused on redesigning Transformer architectures or attention mechanisms. We propose a novel embedding strategy.
 2 related Work
-Transformer-Based Time Series Forecasting.
-Well-designed Transformer-based forecasters have achieved substantial success in time series forecasting tasks.
-They are not limited by the inability of CNN-based methods (Liu et al., 2022) to capture long-term temporal dependencies (Sutskever et al., 2014; Karim et al., 2017), nor do they encounter the vanishing and exploding gradient problems (Bengio et al., 1994; Hochreiter & Schmidhuber, 2019) commonly observed in RNN-based methods (Cho et al., 2014; Lai et al., 2018; Salinas et al., 2020).
-In recent years, advanced models such as Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), and Crossformer (Zhang & Yan, 2023) have primarily introduced novel attention mechanisms and architectural modifications for explicitly modeling long-term dependencies and temporal interactions in time series data.
-Nonetheless, as discussed in the Introduction, Transformer-based forecasters suffer from inherent limitations in effectively modeling time series.
-Refining the attention mechanism and architecture alone cannot address this limitation.
-Unlike prior works, we propose a novel embedding strategy for comprehensive feature extraction on the raw input before it is fed into the attention module, thereby substantially improving the modeling capacity of diverse attention mechanisms.
-As shown in Figure 2.
 Temporal Feature Extraction.
-Previous Transformer-based forecasters, such as FEDformer (Zhou et al., 2022) and MultiPatchFormer (Naghashi et al., 2025), tend to overprioritize temporal feature extraction, introducing various complex mechanisms to model temporal dependencies.
-This design often incurs substantial computational overhead without delivering notable performance gains.
-We argue that the informational content in the temporal dimension is inherently limited and should not be excessively emphasized.
-In contrast, HTMformer performs full temporal feature extraction solely within the embedding layer, thereby significantly reducing computational complexity.
+Previous time series forecasters, especially Transformer-based, such as FEDformer (zhou2022fedformer) and MultiPatchFormer (naghashi2025multipatchformer), tend to overprioritize temporal feature extraction.
+This design often incurs significant computational overhead and can lead to model overfitting.
+In contrast, HTME performs a deep embedding of temporal features at the embedding layer and augments them with multivariate features.
+This process generates semantically richer embedding representations, enabling various models to capture finer-grained features and achieve superior performance.
 Multivariate Correlation.
-iTransformer retains all native Transformer components without modification.
-Instead, it adopts an inverted input representation, enabling the self-attention model to better capture multivariate correlations and thereby yielding satisfactory performance.
-This underscores the importance of modeling multivariate correlations in capturing the semantics of time series.
-In CNN-based predictors, graph convolutional networks (Kipf & Welling, 2017; Yu et al., 2018; Guo et al., 2019) explicitly model correlations among variables by leveraging graph adjacency matrices.
-This paradigm unifies the temporal dynamics of the data with the multivariate structure of variables within a graph-based model, thereby substantially improving performance in time series tasks.
-However, such explicit modeling entails significant computational overhead, particularly in Transformer architectures.
-HTMformer is designed to moderately extract multivariate features and integrate them with temporal information within the embedding layer, yielding semantically enriched embeddings.
-Furthermore, it employs an inverted input strategy to enhance the efficiency and performance of Transformers and reduce the computational costs.
-Figure 2: Previous works have focused on redesigning Transformer architectures or attention mechanisms. We propose a novel embedding strategy.
+GCN (kipf2017gcn; yu2018stgcn; guo2019astgcn) extracts correlations among variables by leveraging graph adjacency matrices, while iTransformer (liu2024itransformer) extracts multivariate correlations using Attention.
+Their success highlights the importance of multivariate correlations.
+However, this approach of explicitly modeling multivariate correlations incurs significant computational overhead.
+In contrast, the HTME strategy employs a lightweight module to independently extract multivariate features without introducing additional complex modules.
+Channel Dependence strategy.
+Recent studies (han2023capacity; li2023revisiting) have found that Channel Dependence model makes it prone to capturing spurious correlations from noise and overemphasizing coincidental patterns in the training data.
+This is because existing models, such as Crossformer, make the multivariate features and temporal features share a single space.
+HTME independently extracts multivariate features and projects them to the temporal feature space, allowing subsequent modules to remain focused on the temporal dimension, which significantly suppresses noise.
 3 HTMformer
 This paper focuses on multivariate time series forecasting tasks, which can be formally defined as follows: Given a historical observation matrix $\mathbf{X}\in\mathbb{R}^{L\times C}$, where $L$ denotes the input sequence length and $C$ represents the number of variables, the goal is to predict the future values $\mathbf{Y}\in\mathbb{R}^{H\times C}$, with $H$ indicating the prediction horizon.
+3.1 HTME Extractor
+To fully capture informative representations from the raw time series, the HTME extractor incorporates two novel components: a temporal feature extractor that can more effectively extract temporal patterns and a lightweight multivariate feature extractor.This parallel design which enable the independent modeling reduces interference between features.
+Finally, we employ a weighted summation for feature fusion, enabling multivariate features to complement temporal features, thereby yielding a richer representation.
+The lightweight design ensures minimal additional computational overhead.
+Within the two branches, we adopting a bottom-up fusion followed by a top-down decomposition strategy.
+This strategy encourages the model to focus on specific dimension, thereby avoiding interference between dimensions, while guiding the model to emphasize import patterns, resulting in more effective feature embeddings.
+The HTME extractor takes raw multivariate time series $X_{\text{in}}\in\mathbb{R}^{L\times N}$ as input and outputs embedding representations that fuse informative features, as illustrated in Figure 3.
+3.1.1 Temporal Feature Extractor
+The key to temporal modeling is to extract both short-term and long-term dependencies.
+It is worth noting that short-term correlations are predominant, while the influence of long-term correlations should not be neglected.
+The strategy of first merging and then progressively decomposing features effectively guides the model to focus step-by-step on the most important patterns.
+First, we adopt a patching strategy to segment short-term temporal representations from the original sequence, and iteratively merge all channels to obtain a short-step representation $T^{N\times(L-K),K}$ .It is worth noting that the process merely partitions the time series without introducing trainable parameters at this stage.
+$P_{1},P_{2},P_{3},\dots,P_{L-K}$ $=X_{1:1+K},X_{2+K},X_{3+K},\dots,X_{L-K:L}$ (1)
+$T^{N\times(L-K),K}$ $=\text{Stack}(P_{1},P_{2},P_{3},\dots,P_{L-K})$ (2)
+Short-term dependencies are particularly important.
+this module first prioritizes the deep extraction of short-term temporal features.
+Specifically, we apply multiple convolutional filters to capture short-term temporal patterns at multiple scales.
+We then fuse the convolutional outputs via a linear layer to obtain a feature representation enriched with short-term dependencies.
+$E^{N\times(L-K),8}$ $=\text{Linear}(\text{Conv}(T^{N\times(L-K),1,K}))$ (3)
+Finally, we unfold the long temporal dimension and employ a second linear layer to capture the long-term dependencies within the sequence, yielding a deep embedding that encapsulates both short-term and long-term temporal dependencies.
+$D_{\text{out}}$ $=\text{Linear}(\text{Flatten}(E^{N\times(L-K),8}))$ (4)
+Compared with existing embedding strategies, our method achieves a more fine-grained modeling of temporal contextual dependencies using only one convolutional layer and two linear layers.
+Figure 3: HTME consists of two independent and parallel feature extraction modules. It employs a strategy of first merging the outputs and then hierarchically decomposing the resulting representation.
+3.1.2 Multivariate Feature Extractor
+We adopt the same merge‑decompose strategy as described above.
+We segment and stack the time series to obtain $T^{N\times(L-K),K}$, following exactly the same procedure as in the aforementioned module.
+Subsequently, we merge the short‑term temporal dimension with the multivariate dimension, and employ a linear layer to process the multi‑step variables, thereby obtaining multivariate features $N^{(L-K),128}$ that account for short‑term temporal lags.
+$N^{(L-K),128}$ $=\text{Linear}(T^{(L-K),K\times N})$ (5)
+Similar to Eq.
+(4), we employ multiple convolutional filters to extract multivariate features at different scales, resulting in a more fine-grained multi-scale multivariate feature block.
+As this module is substantial in size, we therefore apply a specialized GRU network to aggregate the multi-scale features and obtain the integrated representation $B^{N}$.
+$B^{N}$ $=\text{GRU}(\text{Conv}(N^{(L-K),128}))$ (6)
+Considering that the correlations among multivariate variables may differ across time stamps, we apply different convolutional kernels to project the integrated representation $B^{N}$ onto the temporal dimension, thereby obtaining the final representation.
+$V_{\text{out}}$ $=\text{Conv}(B^{N})$ (7)
+here, $D_{\text{out}},V_{\text{out}}\in\mathbb{R}^{N\times D}$.
+We define a learnable fusion weight $\alpha$ to adaptively balance the contributions of the two modules to the downstream model, thereby enhancing the scalability of HTME for adaptation to datasets with diverse characteristics.
+The final output is the sum of the two modules:
+$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$ (8)
+where $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ denotes the embedded representation that fuses temporal and multivariate features.
+HTME is fed into a vanilla Transformer encoder to generate the predictive representations.
 3.2 Structure Overview
-Motivated by the temporal and multivariate correlations in time series, we propose HTMformer, a general-purpose time series forecasting model.
-It adopts the encoder-only architecture and incorporates an HTME extractor to effectively capture temporal and multivariate features.
+We construct HTMformer based on iTransformer, a state-of-the-art and generic Transformer architecture.
+It keeps the native components of the original Transformer intact.
+This implies that other components within HTMformer can be flexibly interchanged with their respective variants.
 As shown in Figure 4, the model includes an HTME extractor, a vanilla Transformer encoder layer, and a projection layer.
-The HTME extractor is designed to deeply extract temporal and multivariate features.
-After the time series is processed, the Transformer encoder layer is used to learn the relationships between multiple channels.
-Finally, a linear model serves as the projection layer.
-RevIN (Kim et al., 2022) is a normalization technique for time series forecasting, which can help models learn and generalize better.
+RevIN (kim2022reversible) is a normalization technique, which can help models learn and generalize better.
 Each data batch is first normalized with RevIN before being fed into the model.
@@
 Finally, a linear model serves as the projection layer to generate the final output $Y\in\mathbb{R}^{H\times C}$.
-3.2 HTME Extractor
-To fully capture informative representations from the raw time series, the HTME extractor incorporates two core components: a temporal feature extractor and a multivariate feature extractor, which enable the independent modeling of two dimensions.
-The final output is obtained by integrating the outputs of these two components.
-The HTME extractor takes multivariate time series as input and outputs embedding representations that fuse informative features, as illustrated in Figure 3 (Left).
-Figure 3: Overall architecture of HTMformer.
-The HTME layer comprises two independent modules for temporal and multivariate feature extraction, directly replacing the original embeddings.
-It also employs inverted input to enable the attention mechanism to capture channels correlations.
-3.2.1 Temporal Feature Extractor
-In the temporal feature extraction module, the time series is first segmented into patches, with each patch treated as a single input token.
-Convolutional operations are then utilized to capture short-term temporal patterns within local receptive fields.
-The convolutional outputs are flattened into a high-dimensional tensor that integrates richer temporal information.
-Then, a linear projection layer is applied to further integrate local temporal features.
-The extracted features are subsequently projected along the temporal dimension to model the long-term correlations, and the tensor is reshaped to generate the final representation.
-Compared with conventional embedding layers, combining convolution with patch-based tokenization enables more accurate modeling of temporal contextual dependencies.
-By treating each channel as an independent input, interference from other features is reduced, thereby enhancing the model’s capability to capture temporal features effectively.
-$N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}$ $=\operatorname{Conv}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (2)
-$E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}$ $=\operatorname{Linear}\bigl(\operatorname{Flatten}\bigl(N_{1}^{4\times k},\,N_{2}^{4\times k},\,\ldots,\,N_{(L-k)N}^{4\times k}\bigr)\bigr),$ (3)
-$D_{\text{out}}$ $=\operatorname{Linear}\bigl(E_{1}^{\text{dim}},\,E_{2}^{\text{dim}},\,\ldots,\,E_{N}^{\text{dim}}\bigr).$ (4)
-3.2.2 Multivariate Feature Extractor
-In the multivariate feature extraction module, the time series is first partitioned into patches to accommodate potential temporal lags.
-The features of each patch are then flattened to aggregate local information, and a linear layer is applied to model the correlations among input variates.
-These learned relationships are aggregated across time through temporal convolution.
-The resulting output is fed into a GRU (Cho et al., 2014) network, whose gating mechanism selectively emphasizes or suppresses different inputs, thereby adaptively modulating the contribution of historical information.
-Finally, the features fused with multivariate information are expanded along the temporal dimension via a convolution layer to match the output of the temporal module.
-$\left(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\right)$ $=\operatorname{Flatten}\bigl(\operatorname{Patch}(X_{\text{in}})\bigr),$ (5)
-$M^{N}$ $=\operatorname{GRU}\bigl(\operatorname{Linear}\bigl(P_{1}^{C_{k}},P_{2}^{C_{k}},\ldots,P_{L-k}^{C_{k}}\bigr)\bigr),$ (6)
-$V_{\text{out}}$ $=\operatorname{Conv}\bigl(M^{N}\bigr),$ (7)
-here, $D_{\text{out}},V_{\text{out}}\in\mathbb{R}^{N\times D}$.
-We define a learnable fusion weight $\alpha$ to adaptively balance the contributions of the two modules to the downstream model, thereby enhancing the scalability of HTME for adaptation to datasets with diverse characteristics.
-The final output is the sum of the two modules:
-$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$ (8)
-where $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ denotes the embedded representation that fuses temporal and multivariate features.
-HTME is fed into a vanilla Transformer encoder to generate the predictive representations.
+Figure 4: Architecture of HTMformer. The HTME directly replacing the iTransformer embeddings. It also employs inverted input to enable the attention mechanism to capture channels correlations.
 3.3 Encoder and Projection
 The Transformer encoder layer is composed of a self-attention module and a position-wise feed-forward network.
-As temporal features have been fully extracted in the embedding stage, we adopt an inverted input design, which enables the self-attention mechanism to directly model global channel correlations.
+We adopt an inverted input design, which enables the self-attention mechanism to directly model global channel correlations.
 This inverted structure avoids forced alignment of variables within timestamps and facilitates the concatenation of temporal and multivariate dimensions.
-Consequently, it enables the attention mechanism to learn more appropriate sequence representations, leading to more accurate predictions.
-This reduces computational complexity from $\mathcal{O}(L^{2})$ in traditional methods to $\mathcal{O}(N^{2})$.
+Consequently, it enables the attention mechanism to learn more appropriate sequence representations.
 The attention mechanism projects $Y_{\text{in}}\in\mathbb{R}^{N\times D}$ to generate query matrix $Q_{k}$, key matrix $K_{k}$, and value matrix $V_{k}$, where $d_{k}$ denotes the projection dimension for each attention head.
@@
 $\left({O}_{k}\right)^{T}$ $=\text{Softmax}\left(\frac{Q_{k}K_{k}^{T}}{\sqrt{d_{k}}}\right)V_{k}.$ (10)
-The self-attention block additionally consists of layer normalization and a position-wise feed-forward network with residual connections.
-It enables the model to learn cross-channel relationships and yield refined representations for each channel.
-iTransformer(Liu et al., 2024b) observes that, for numerical time-series forecasting tasks, a linear layer is sufficient to generate predictions, since forecasting can be realized by mapping variate-level representations without resorting to complex decoder architectures.
+The self-attention block additionally consists of layer normalization and a position-wise feed-forward network.
 Finally, a linear projection head serves as the decoder, mapping the learned representations to the target dimensionality and yielding the final predictions.
-Unlike other models that deeply mine multidimensional correlations (Ying et al., 2021; Lai et al., 2018), our model does not explicitly model the associations between variables.
-Within HTME, a novel strategy is adopted to obtain high-dimensional representations that integrate multivariate information, which are subsequently mapped onto the temporal dimension.
-We utilize multivariate features as a complement to temporal features to learn richly informative embeddings, thereby achieving the goal of deeply extracting the hybrid temporal and multivariate characteristics of time series.
+iTransformer (liu2024itransformer) observes that this Transformer‑based predictor architecture is more efficient for time series forecasting tasks.
 3.4 Complexity Analysis
-HTMformer is composed of two primary components: the HTME extraction module and a Transformer encoder.
-The model’s computational complexity is primarily dictated by the Transformer encoder, as the introduction of the inverted input design results in both time and memory complexity scaling as $\mathcal{O}(N^{2})$.
-Compared with the state-of-the-art Transformer forecaster PatchTST (Nie et al., 2023) and the vanilla Transformer, HTMformer attains lower complexity because the number of variate tokens is fixed and independent of the lookback length, which confers clear advantages for long-term time series forecasting.
-Moreover, as the key component, the HTME extraction module is flexibly compatible with Transformer variants integrating efficient attention mechanisms, such as Informer or Reformer, It allows for further reductions in computational and memory costs and broadens the model’s applicability to a wider range of forecasting scenarios (see Appendix C).
+HTME consists of two parallel extraction modules.
+The time and space complexities of the Feature Extraction Modules are both O(LN).
+For each dimension, the computational complexity is linear, which guarantees the high efficiency and scalability of HTME.
+The HTME module does not introduce significant additional computational overhead.
+The primary computational cost of HTMProdectors stems from the underlying base predictor they are built upon.
 4 Experiments
-To evaluate both the effectiveness of the proposed HTME strategy in enhancing Transformer-based forecasters and its capacity for scalability, we apply HTME to various Transformer variants.
-Notably, HTME consistently and significantly improves the performance of Transformers.
-Moreover, we conduct a comprehensive evaluation of the proposed HTMformer on diverse time series forecasting and analysis tasks to further validate the advantages of the HTME strategy.
-Across various key tasks, HTMformer consistently outperforms current state-of-the-art methods.
+To validate the effectiveness and transferability of the HTME strategy, we integrated it into diverse forecasting backbones and evaluated its impact on predictive performance.
+Moreover, we conducted a comprehensive evaluation of the proposed HTMformer on diverse forecasting tasks to further validated the advantages of the HTME strategy.
+We also evaluated the computational overhead of HTME, showing that it introduces only modest additional cost.
 The detailed experimental configurations and implementation details can be found in Appendix D.
 4.1 Experiment Settings
-All models are evaluated under identical settings, following the Time-Series-Library (Wang et al., 2024b), to ensure the fairness of comparison and the reliability of the drawn conclusions.
-Datasets.
-We evaluated the performance of HTMformer on eight widely used datasets, including Weather, Traffic, Electricity, ETTh2 adopted in Autoformer (Wu et al., 2021), Solar-Energy proposed in LSTNet (Lai et al., 2018), and three PEMS datasets (PEMS03, PEMS04, PEMS08) investigated in SCINet (Liu et al., 2022).
+All models are evaluated under identical settings, following the Time-Series-Library (wang2024tssurvey), to ensure a fair comparison and reliable conclusions..
+We evaluated the performance of HTMformer on eight widely used datasets, including Weather, Traffic, Electricity, ETTh2 adopted in Autoformer (wu2021autoformer), Solar-Energy proposed in LSTNet (lai2018modeling), and three PEMS datasets (PEMS03, PEMS04, PEMS08) investigated in SCINet (liu2022scinet).
 These datasets span diverse domains and sampling frequencies, providing a comprehensive experimental scenario for model evaluation.
-Baselines and metrics.
-We employ eight carefully selected state-of-the-art Transformer-based models and widely recognized non-Transformer-based models in the time series forecasting domain as baselines, encompassing iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Dlinear (Zeng et al., 2023), WPMixer (Murad et al., 2025), MultiPatchFormer (Naghashi et al., 2025), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023).
-Mean Squared Error (MSE) and Mean Absolute Error (MAE) are utilized as the assessment criteria.
+We employ eight carefully selected state-of-the-art models from three different architectures.
 4.2 Performance Promotion with HTME
-To assess the effectiveness and scalability of HTME, it is integrated into the Transformer (Vaswani et al., 2017) and its variants, as shown in Table 1.
-Specifically, HTME achieves consistent improvements across different Transformer variants, with average performance gains of 35.8%, 34.3%, 43.6%, and 31.9% for Transformer, Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022), respectively.
-These results demonstrate that HTME captures more informative time series representations and significantly strengthens the modeling capacity of Transformer-based forecasters.
-More importantly, the temporal feature extraction module and the multivariate feature extraction module within the HTME layer operate as entirely decoupled components.
-Consequently, the multivariate representations produced by HTME can be seamlessly integrated with other embedding schemes that emphasize temporal characteristics, thereby further enhancing the applicability and versatility of HTME.
-The HTME framework is broadly applicable to Transformer-based forecasting architectures, allowing for flexible utilization of various attention mechanisms tailored to different application scenarios.
+To assess the effectiveness and scalability of HTME, the vanilla self-attention mechanism (vaswani2017attention) and its variants are integrtated into the HTMformer framework, as shown in Table 1.
+Specifically, HTMformers achieve consistent improvements across different variants, with average performance gains of 35.8%, 34.6%, 43.6%, 31.9%,and 33.1% for Transformer, Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention), respectively.
 Table 1: Performance improvementst by HTME. Full results can be found in Appendix E.1.
-Models Transformer Reformer Informer Flowformer Flashformer
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Electricity Original
+Models Transformer Reformer Informer Flowformer Flashformer iTransformer DLinear NLinear SegRNN
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ECL Original
 0.271
@@
 0.364
+0.244
+0.290
+0.225
+0.318
+0.217
+0.295
+0.194
+0.288
 +HTME
@@
 0.287
+0.194
+0.288
+0.219
+0.314
+0.208
+0.291
+0.190
+0.283
 Promotion
@@
 21.1%
+20.4%
+0.6%
+2.6%
+1.2%
+4.1%
+1.3%
+2.0%
+1.7%
 Weather Original
@@
 0.559
+0.259
+0.279
+0.269
+0.290
+0.274
+0.299
+0.254
+0.301
 +HTME
@@
 0.280
+0.252
+0.267
+0.261
+0.286
+0.261
+0.290
+0.249
+0.296
 Promotion
@@
 49.9%
+2.7%
+4.0%
+2.9%
+1.3%
+4.7%
+3.0%
+1.9%
+1.6%
 Traffic Original
@@
 0.362
+0.474
+0.317
+0.672
+0.418
+0.632
+0.391
+0.761
+0.397
 +HTME
@@
 0.319
+0.473
+0.321
+0.609
+0.383
+0.611
+0.384
+0.750
+0.392
 Promotion
@@
 11.8%
+0.2%
+0.0%
+9.3%
+8.1%
+3.3%
+1.7%
+1.4%
+1.3%
 Promotion Avg
@@
 33.1%
+4.6%
+4.2%
+3.0%
+1.7%
+To demonstrate the strong scalability of HTME strategy, we apply it to diverse architectures, including the Transformer-based iTransformer, the linear-based DLinear, NLinear and the RNN-based SegRNN.
+For these models, we supplement the original input with multivariate features.
+The results show that HTME consistently improves the forecasting performance across all these models.
+The HTME strategy consistently boosts the performance of a wide range of forecasting models.
+More importantly, the multivariate representations produced by HTME can be seamlessly integrated with other embedding schemes that emphasize temporal characteristics.
+These further enhance the effectiveness and scalability of HTME.
 4.3 Long-term Forecasting
-Long-term forecasting holds significant importance in various application domains, including meteorology, traffic management, and energy utilization.
-We conduct a comprehensive evaluation of the effectiveness of HTMformer in long-term forecasting.
+Long-term forecasting is critical for real-world applications such as meteorology, traffic management, and energy systems.
 Table 2 presents the experimental forecasting results.
 Within each row, the lowest averaged MSE and MAE values across four prediction horizons are highlighted in red, and the second-lowest values are underscored in blue.
-HTMformer outperforms other models in long-term forecasting across various datasets, especially on high-dimensional datasets such as Traffic.
-We further provide a visualization of the long-term forecasting results produced by HTMformer (see Appendix I).
-HTMformer effectively captures the periodic characteristics of time series, delivers accurate trend forecasts for complex sequences, and demonstrates strong robustness against large-scale fluctuations.
-In conclusion, HTMformer demonstrates significant performance improvements and tangible practical benefits in long-term forecasting tasks.
-Table 2: Average performance for long-term forecasting over prediction horizons $H\in\{96,192,336,720\}$ with fixed lookback $L=96$.
-Notably, MPFormer is short for MultiPatchFormer.
-Full results are listed in Appendix E.2.
+HTMformer achieves the best average performance on most datasets, with particularly strong gains on high-dimensional benchmarks such as Traffic.
+We further provide qualitative visualizations of the predicted trajectories (see Appendix J).
+Overall, HTMformer delivers significant improvements in long-term forecasting, highlighting its practical value.
+Table 2: Average performance for long-term forecasting over prediction horizons $H\in\{96,192,336,720\}$. MPFormer is MultiPatchFormer. Full results are listed in Appendix E.2.
 Models
@@
 SegRNN
+nie2023time
+zeng2023dlinear
+zhou2022fedformer
+naghashi2025multipatchformer
+murad2024wp
+wang2023timemixer
+lin2023segrnn
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 4.4 Short-term Forecasting
-Short-term time series forecasting tasks are also prevalent across a wide range of application domains.
+Short-term time series forecasting tasks is also prevalent in many real-world applications.
 Table 3 presents the experimental results on the PEMS (PEMS03, PEMS04, and PEMS08) datasets.
-The proposed HTMformer demonstrates better performance across all three PEMS datasets.
-It outperforms the second-best model, MultiPatchFormer, yielding a 21.7% reduction in MSE and a 12.0% reduction in MAE, respectively.
+The proposed HTMformer demonstrates BEST performance across all three PEMS datasets, outperforming the second-best model, MultiPatchFormer, with a 21.7% reduction in MSE and a 12.0% reduction in MAE.
 We attribute this performance gain to the exploitation of rich multivariate representations in HTME.
 Since short-term forecasting, constrained by the limited prediction length, does not entail intricate long temporal dependencies, it renders inter-variable correlations more salient.
-These results validate the robust performance of the HTMformer on short-term forecasting tasks, further complementing its strong performance in long-term forecasting.
+These results further demonstrate HTMformer’s strong and robust performance, complementing its long-term results.
 Table 3: Short-term forecasting results, averaged across four prediction lengths $H\in\{12,24,48,96\}$ with a fixed lookback window of $L=96$. Full results are available in Appendix E.3.
@@
 TimeMixer
+nie2023time
+zeng2023dlinear
+zhou2022fedformer
+naghashi2025multipatchformer
+murad2024wp
+wang2023timemixer
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 0
-Compared with iTransformer (Liu et al., 2024b), which explicitly models multivariate correlations, and PatchTST (Nie et al., 2023), which excels at capturing fine-grained temporal patterns, HTMformer achieves broadly superior performance on both long-term and short-term forecasting tasks.
-Even when compared with the previously best-performing Transformer-based forecaster, MultiPatchFormer, HTMformer consistently achieves either superior or comparable predictive accuracy.
-This performance gain is particularly pronounced on high-dimensional time series datasets, such as Traffic and Solar-Energy.
-Such effectiveness is attributed to the use of the HTME extractor, which facilitates multi-dimensional feature fusion in time series.
-HTME adaptively integrates temporal and multivariate representations to achieve comprehensive, multidimensional feature complementarity.
-Furthermore, it encapsulates richer and more discriminative semantic representations, enabling Transformer-based forecasters to more fully leverage sequence information.
-This demonstrates the broad applicability across diverse datasets, thereby highlighting the effectiveness of HTME.
 5 Model Analyses
 5.1 Ablation Studies
-To evaluate the contribution and reliability of each module within HTMformer, we conducted ablation studies (see Appendix F).
-Specifically, we examined two different variants of HTMformer: 1) HTMformerV1, which utilizes only the temporal feature extraction module of the HTME extractor; 2) HTMformerV2, which utilizes only the multivariate feature extraction module of the HTME extractor.
-We also examined another variant 3) iTransformerV3 (iTransV3), which integrates the multivariate feature extraction module of the HTME extractor into the embedding layer of iTransformer.
-We adopt iTransformer (Liu et al., 2024b) as the state-of-the-art baseline.
-Sharing a highly similar architectural design with HTMformer, both models employ a dedicated embedding module and process inverted inputs before feeding them into the Transformer encoder.
-However, in iTransformer, the embedding layer conducts feature extraction exclusively along the temporal dimension.
-Notably, since all experiments were performed under identical configurations, the HTMformer variants are directly comparable to the other models.
-We also adopt MultiPatchFormer (Naghashi et al., 2025), attaining the second-best performance in various forecasting tasks, as a representative proxy for other forecasting models.
-Specific results are documented in Table 4.
+To evaluate the contribution of each component in HTME, we conduct ablation studies (see Appendix F).
+Specifically, we examined three different variants of HTMProductors: 1) Productor, the baseline forecasting model.
+2) ProductorV1, which utilizes only the multivariate feature extraction module of the HTME extractor.
+3) ProductorV2, which utilizes the multivariate feature extraction module of the HTME extractor to supplement the embedding representations.
+We further introduce two variants: HTMformer, builds upon iTransformer by incorporating the complete HTME extractor; HTMformerV1, replaces the original embedding layer in iTransformer with the temporal feature extractor from HTME.
+We adopt iTransformer (liu2024itransformer) as the state-of-the-art baseline.
+The results are summarized in Table 4.
 Table 4: Results (Average) of different variants under prediction lengths $H\in\{96,192,336,720\}$. The input lookback $L$ is set to 96.
@@
 HTMformer
-HTMformerV1
-HTMformerV2
+HTMformer-A
+iTransV2
 iTransformer
-iTransV3
+iTransV1
 MPFormer
@@
 Effect of the temporal extractor.
-Compared with other time series forecasting models, HTMformerV1 already outperforms several baselines and suffers only a marginal accuracy drop even when benchmarked against state-of-the-art approaches.
-Blue numbers indicate the cases where HTMformerV1 outperforms iTransformer.
-This observation suggests that our temporal feature extraction module has effectively captured informative temporal patterns.
-Effect of the multivariate features.
-To facilitate an intuitive comparison, we evaluate HTMformerV1 against HTMformerV2, with numerical values underlined when HTMformerV2 achieves better performance over HTMformerV1.
-Across most datasets, particularly those with a large number of variables, HTMformerV2 performs substantially worse than other time series forecasting models, highlighting that effective forecasting primarily depends on accurate temporal dependency modeling.
-Nevertheless, on certain datasets, HTMformerV2 attains superior or comparable results relative to most baseline models.
-This emphasizes that multivariate features are crucial and indispensable sources of information in time series forecasting.
+The experimental results show that HTMformer-A and iTransformer outperform iTransformerV1 in most cases.
+This suggests that the primary features of time series are embedded within the temporal dimension.
+Furthermore, HTMformer-A demonstrates superior performance over iTransformer across the majority of which showwd in blue.
+This indicates that deeply mining temporal patterns can further enhance prediction performance.
+These findings underscore the dominant role of temporal features in time series forecasting.
 Effect of the hybrid strategy.
-We further conduct comparative evaluations between HTMformer and HTMformerV1, as well as between iTransformerV1 and the original iTransformer.
+On the PEMS dataset, which exhibits strong multivariate correlations, PredictorV2 achieved superior performance.
+This indicates that multivariate correlation is a significant pattern in prediction.
+However, for most datasets, multivariate feature extraction can introduce substantial noise.
+Effect of the hybrid strategy.
+We further conduct comparative evaluations between Predictors and PredictorV1.
 Red numbers indicate cases where the versions augmented with multivariate integration in the embedding layer outperform their original counterparts.
-Across most datasets, both HTMformer and iTransformerV1 consistently achieve superior performance.
+Across most datasets, PredictorV1 consistently achieved superior performance.
 This provides strong evidence for the effectiveness of our proposed strategy for incorporating multivariate features into embedding representations.
+The multivariate features consistently enriched the temporal feature embeddings.
 5.2 Model Efficiency
-To assess the computational efficiency of HTMformer, we perform a comparative analysis of its training time, GPU memory footprint, and parameter count against state-of-the-art models, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023).
-All models are evaluated on the same GPU under identical lookback windows, prediction horizons, and batch sizes to ensure fair comparison.
-As shown in Figure 4, using the state-of-the-art MultiPatchFormer as the reference baseline, HTMformer achieves a training runtime approximately one-third of that of MultiPatchFormer, while requiring only 20% to 45% of its GPU memory footprint.
+To assess the computational efficiency of HTME, we compare the floating-point operations (FLOPs) and the number of parameters of iTransformer, DLinear, RLinear, and SegRNN against their HTME versions.
+Owing to its carefully designed architecture, the computational overhead introduced by the HTME multivariate feature extraction module is nearly constant: for complex Transformer-based models it adds negligible extra cost, and for simple models the overall overhead remains modest.
+We perform a comparative analysis of HTMformer against MultiPatchFormer (naghashi2025multipatchformer), and PatchTST (nie2023time).
+As shown in Figure 5, taking the state-of-the-art MultiPatchFormer as the reference baseline, HTMformer reduces training runtime to roughly one third while using only 20%–45% of the GPU memory.
 Moreover, HTMformer’s model parameters are just half those of MultiPatchFormer.
 Importantly, these gains are achieved without sacrificing predictive accuracy, as HTMformer consistently attains superior or comparable MSE and MAE across eight benchmark datasets.
-These results indicate that HTMformer enables faster training and inference under the same hardware conditions, and its smaller memory footprint also enhances the model’s applicability (see Appendix G).
+The HTME strategy enables simple backbones to achieve performance comparable to much more complex architectures by introducing only a lightweight module (see Appendix G).
 Figure 5: We compare HTMformer with three state-of-the-art models on eight benchmark datasets. The input sequence length is uniformly set to 96, and the prediction sequence length is 192.
-It is noteworthy that the ablation studies and the long-term forecasting experiments were conducted under identical configurations, rendering the results directly comparable.
-We observe that for time series data, using patching and convolutions for local patterns and linear layers for long-range dependencies is generally sufficient for effective temporal modeling.
-Excessive modeling of temporal dependencies yields limited gains in forecasting accuracy while substantially increasing compute and memory overhead.
-In this study, we adopt a hybrid design in the embedding layer, employing two lightweight models to extract multidimensional features.
-The resulting hybrid embeddings extract richer informative representations from the sequences, thereby assisting Transformer-based forecasters in modeling time series.
-This approach not only improves predictive performance but also markedly reduces both computational demand and memory footprint.
+5.3 Analysis of HTME
+To better understand the feature space induced by HTME, we visualize the parameters defined in Equation (6), which maps multivariate feature embeddings onto the temporal feature space.
+The visualization shows that temporal features are consistently assigned substantially larger weights than multivariate features.
+The vast majority of features in time series are contained within the temporal dimension (hyndman2021forecasting).
+The multivariate dimension also contains a number of features (Pravilovic2013SpatialForecasting).
+However, it does not warrant the use of an additional complex module for its modeling.
+In the visualization, temporal features exhibit a line-shaped distribution, indicating a consistent temporal correlation within the series.
+In contrast, multivariate features appear as scattered point-like clusters, emphasizing a few salient variables.
+With this design, HTME augments temporal embeddings with multivariate signals: the embeddings fed into the main backbone are still treated as temporal features, but they are now enriched with additional semantic information, thereby enhancing the backbone’s modeling capacity.
+This approach avoids introducing extra modules into the backbone network and ensures the consistency of the feature space (see Appendix H).
 6 Conclusion and Limitations
-Motivated by the inherent properties of multivariate time series, we introduce the HTME strategy.
+In this work, we propose the HTME strategy.
 The HTME module jointly encodes temporal dynamics and inter-variate dependencies present in the raw sequences, yielding embedding representations that are richer and more informative than those produced by prior designs.
-By inverting the input representation, the attention mechanism is empowered to more effectively model inter-channel dependencies.
-Experimental results demonstrate that HTMformer delivers state-of-the-art forecasting performance with faster training and a reduced GPU memory footprint, thereby validating the effectiveness of injecting multivariate information into the embedding space.
-Moreover, the proposed HTME module in HTMformer is seamlessly compatible with diverse Transformer architectures, underscoring its strong scalability.
-Through numerous experiments, this paper further highlights that incorporating both temporal and multivariate information enables Transformer-based predictors to make more accurate predictions.
-The detailed discussion is provided in Appendix J.
+This allows the predictor to avoid introducing additional modules for modeling multivariate correlations.
+Consequently, HTME consistently enhances the modeling capabilities of various models, enabling lightweight models to achieve performance comparable to that of complex ones.
 However, time series data exhibit complex correlations, and simply adding multivariate features to the temporal dimension through the embedding layer cannot fully model such a complex pattern.
 Jointly modeling spatiotemporal dependencies in time series remains an open problem that requires further exploration.
-Future improvements in Transformers can focus on devising methods to efficiently capture multivariate correlations.
 7 AI Usage Statement
@@
 10 Acknowledgements
-Appendix A Experiments of Different Embedding Methods
-To demonstrate that Transformer-based forecasters have limited ability to understand sequences and the necessity of deeply mining sequence features in the embedding layer, we compared the prediction performance of Transformer (Vaswani et al., 2017), iTransformer (Liu et al., 2024b), and PatchTST (Nie et al., 2023) with different embedding approaches on four datasets: ECL, Weather, ETTh2, and Solar.
+The appendix is divided into several sections, each giving extra information and details.
+Appendix A Discussion on Multivariate Feature Extraction
+Recent studies (han2023capacity; li2023revisiting) have found that Channel Dependence ideally gains from higher capacity, while Channel Independence can significantly enhance performance due to sample scarcity.
+As most current forecasting benchmarks are not sufficiently large, the complex design of the Channel Dependence model makes it prone to capturing spurious correlations from noise and overemphasizing coincidental patterns in the training data.
+In time series forecasting, an increasing number of models employ Channel Independence (nie2023time) or exclusively focus on modeling the temporal dimension.
+These models achieved good performance.
+However, they inevitably overlook the multivariate correlations, resulting in the loss of some information.
+We believe that multivariate correlations should not be ignored.
+Past hybrid models explicitly model multivariate correlations by incorporating various extra and complex modules within their backbone networks.
+This strategy mixes temporal and multivariate features within the same space.
+However, constrained by the sample size, the multivariate extraction process can introduce significant noise and spurious patterns.
+Moreover, the additional multivariate features dilute the temporal features that should be dominant, leading to a disorganized feature space.
+Moreover, the additional and complex module for multivariate feature extraction introduces significant computational overhead.
+Appendix B Experiments of Different Embedding Methods
+To demonstrate that Transformer-based forecasters have limited ability to understand sequences and the necessity of deeply mining sequence features in the embedding layer, we compared the prediction performance of Transformer (vaswani2017attention), iTransformer (liu2024itransformer), and PatchTST (nie2023time) with different embedding approaches on four datasets: ECL, Weather, ETTh2, and Solar.
 The input sequence length is set to 48, 96, 144, 192, 240, and 288, with the prediction length being 96.
 The experimental results are shown in Figure 6.
-Experimental results reveal that increasing the input sequence length initially leads to a significant decline in prediction loss.
+Experimental results reveal that employing more sophisticated embedding methods to capture complex spatiotemporal dependencies can effectively enhance the performance of Transformers.
+These findings suggest that Transformer-based models exhibit limitations in proficiency in capturing effective features.
+This limitation arises because the embedding layer fails to effectively represent the complex characteristics of time series.
+Constructing a sophisticated embedding method to extract semantically richer representations from the raw sequence can significantly enhance the performance of Transformers.
+We also noticed that increasing the input sequence length initially leads to a significant decline in prediction loss.
 However, as the sequence length continues to grow, this benefit markedly lessens, and in some cases, prediction loss may even increase.
-These findings suggest that Transformer-based models exhibit limitations in proficiency in capturing effective features from raw series data.
 Simply increasing the lookback window does not consistently capture more informative patterns and will also incur additional computational overhead.
-This is because time series forecasting focuses on capturing short-term dependencies (Kim et al., 2025) and the temporal dimension inherently offers limited contextual information.
-Therefore, traditional Transformer-based forecasters encounter bottlenecks in reliably improving prediction accuracy.
-Notably, employing more sophisticated embedding methods to capture complex spatiotemporal dependencies can effectively address this problem.
-Motivated by these findings, we propose a novel embedding strategy that moderately extracts features along the temporal dimension to balance prediction accuracy and computational complexity, while integrating multivariate features to enhance informative content, thereby enabling the efficient acquisition of high-quality embeddings.
-Combined with input inversion, this approach brings comprehensive improvements to Transformer-based forecasters.
+This is because time series forecasting focuses on capturing short-term dependencies (kim2025comprehensivesurveydeeplearning) and the temporal dimension inherently offers limited contextual information.
+Therefore, we argue that it is highly beneficial to consider multivariate correlations during the embedding stage.
+Doing so can greatly enrich the information content of the embedded representations and, in turn, enhance the modeling capability of subsequent modules.
+Motivated by these findings, we propose a novel strategy that moderately extracts features along the multivariate dimension to enhance informative content, thereby enabling the efficient acquisition of high-quality embeddings.
 Figure 6: On the ECL, Weather, ETTh2, and Solar, we compared three models with input lengths of 48, 96, 144, 192, 240, 288 and an output length of 96. The evaluation metrics were MSE and MAE.
-Appendix B Limitations of Conventional Embedding
-Time series data inherently exhibits both temporal and multivariate (spatial) characteristics.
-A growing body of research has demonstrated the effectiveness of explicitly capturing intra- and inter-channel dependencies (Zhang & Yan, 2023; Liu et al., 2024b; a).
-In prevailing Transformer-based forecasting frameworks, the embedding stage typically consists of three main components: timestamp embedding, positional embedding, and time series convolution.
-These components are primarily designed to capture temporal patterns while often overlooking the intrinsic correlations among different variables within the data.
-In fact, many types of time series data exhibit significant multivariate correlations, such as traffic flow and weather forecasting (Li et al., 2018; Wu et al., 2020b), where sensors are installed at fixed spatial locations.
-However, conventional embeddings capture only temporal patterns, while the subsequent attention layer also primarily models temporal dependencies, thereby overlooking multivariate dependencies and leading to information loss.
-Due to the overemphasis on temporal modeling and the loss of multivariate information under such designs, existing Transformer-based forecasters have encountered significant bottlenecks in capturing the contextual relationships essential for accurate multivariate time series forecasting.
-Furthermore, the temporal feature extraction approach of the embedding layer, inspired by large language models (LLMs), is ill-suited for time series analysis.
-Unlike the Transformers used in natural language and vision tasks (Kirillov et al., 2023; OpenAI, 2023), which require learning dependencies among thousands to millions of tokens, Transformers in time series forecasting generally operate within relatively limited contexts, involving no more than hundreds of time series tokens.
-A smaller number of tokens makes it difficult for the attention mechanism to learn clear temporal correlations.
-However, time series convolution is limited to extracting local patterns from the original sequence and lacks the capability for fine-grained feature extraction.
-Additionally, PRformer (Yu et al., 2025) highlights that traditional Transformer algorithms attempt to learn sequence variation patterns by combining positional encoding with multivariate data at different time instants.
-However, real-world time series are frequently non-stationary, and the assumption of translation invariance makes it difficult for the Transformer’s self-attention mechanism to learn correct temporal dependencies.
-Our experimental results (see Appendix A) indicate that the quality of the embeddings has a substantial impact on the performance of Transformer-based forecasters.
-Therefore, it is imperative to redesign the embedding model to capture more effective representations.
-The HTME extractor discards positional encoding and treats timestamp information as a separate variable, concatenated to the time series inputs.
-This method has enhanced the model’s ability to learn the relationship between the sequence and time.
-It employs patch-based operations and convolution to more effectively capture short-term dependencies, and a linear layer to capture long-term correlations.
-This method can extract temporal features more effectively and is better suited for Transformer-based architectures in time series forecasting tasks.
-We further employ a well-designed multivariate feature extractor to enrich the embedding representations with multivariate features.
-Notably, we do not explicitly model multivariate correlations, as doing so would incur substantial computational overhead.
-Finally, it incorporates multivariate features into temporal features to complement and adaptively balances the contributions of the two feature components, thereby enhancing the scalability of the embeddings.
-The resulting HTME incorporates rich temporal information and thoroughly accounts for multivariate correlation effects.
-It is more effective than existing embeddings in extracting informative representations from raw time series.
-This approach has already achieved good results in temporal feature extraction, making it unnecessary to overemphasize temporal correlations through the attention mechanism.
-So HTMformer utilizes inverted input to enable the attention mechanism to capture correlations among multiple features.
-iTransformer (Liu et al., 2024b) demonstrates that this approach leads to improved performance.
-Appendix C Time and Memory Complexity
-Table 5 provides a detailed comparison of the time complexity and memory usage across various Transformer-based forecasters.
-$L$ denotes the input sequence length, $N$ denotes the number of variables, and $S$ denotes the stride length of PatchTST.
-Table 5: Time and memory complexity of different Transformer-based forecasters.
-Methods Time Memory
-Transformer $O(L^{2})$ $O(L^{2})$
-Informer $O(L\log L)$ $O(L\log L)$
-Reformer $O(L\log L)$ $O(L\log L)$
-PatchTST $O({L^{2}}/{S^{2}})$ $O({L^{2}}/{S^{2}})$
-iTransformer $O(N^{2})$ $O(N^{2})$
-HTMformer $O(N^{2})$ $O(N^{2})$
+Appendix C Detail of HTME
+To demonstrate the correctness of the HTME extractor, we provide a detailed theoretical explanation for each of its steps.
+The HTME temporal feature extraction module consists of four steps: Patching, Stacking, short-term feature extraction, and long-term feature extraction.
+The purpose of the Patching step is to segment the input into short-term temporal sequences.
+The Stacking step is inspired by the concept of channel independence, which has been proven to be more effective (see Appendix A).
+Time series data inherently exhibit strong short-term correlations and weaker long-term dependencies (see Appendix C).
+Motivated by this, we first employ convolutional layers to extract multi-scale representations from the short-term sequences, followed by a linear layer to capture local features.
+Subsequently, another linear layer is used to model the long-term dependencies among these features.
+This hierarchical process yields a more informative temporal feature representation.
+The key to multivariate feature extraction is to reduce overfitting and mitigate noise (see Appendix A).
+For the multivariate feature extraction module, we adopt the exact same logic and methodology.
+It is worth noting that we cannot simply treat the variables as being independent of the temporal dimension.
+We concatenate the multivariate data over short time windows and then use a linear layer to extract features, to account for the effect of time delays.
+A change occurring at one node may exert influence on other nodes at the subsequent time step.
+This extraction strategy, which fuses short timesteps, enables the model to deeply consider the complex multivariate correlations within time series, thereby effectively suppressing over-attention to coincidental patterns.
+$N^{(L-K),128}$ $=\text{Linear}(T^{(L-K),K\times N})$
+Subsequently, we employ a limited number of convolutional layers to extract multivariate features at different scales.
+Fewer convolutional kernels allows the model to focus on significant patterns, thereby substantially mitigating overfitting.
+Then, a specialized linear layer that GRU network integrated multi-scale features to further extract effective features.
+$B^{N}$ $=\text{GRU}(\text{Conv}(N^{(L-K),128}))$
+Finally, small convolutional kernels are used to fine-tune the multivariate features along the temporal dimension, which prevents the model from learning fixed patterns.
+$V_{\text{out}}$ $=\text{Conv}(B^{N})$
+Finally, we employ a parameter to adjust the weights of the two spaces.
+This parameter is initialized with a large value to guide subsequent modules to focus primarily on temporal features.
+During training, HTME adaptively considers the influence of multivariate features to generate high-quality embedding representations.
+$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$
 Appendix D Detailed Experiment Settings
@@
   • PEMS. Includes PEMS03, PEMS04, and PEMS08 datasets, each containing public transportation network data collected at 5-minute intervals in California.
-To ensure the consistency of our experimental approach and make sure there are no data leakage issues, we split the ETT and PEMS datasets into the training, validation, and test sets, with a 6:2:2 ratio and other datasets with a 7:1:2 ratio, following the protocols recommended in recent literature (Wu et al., 2021; Zeng et al., 2023).
+To ensure the consistency of our experimental approach and make sure there are no data leakage issues, we split the ETT and PEMS datasets into the training, validation, and test sets, with a 6:2:2 ratio and other datasets with a 7:1:2 ratio, following the protocols recommended in recent literature (wu2021autoformer; zeng2023dlinear).
 In all forecasting experiments, the lookback window was fixed at 96 time steps, with prediction horizons of 96, 192, 336, 720 for the long-term experiments and 12, 24, 48, 96 for the short-term experiments.
@@
 Transportation
-  • $\ast$ The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (Goerg, 2013). A larger value indicates better predictability.
+  • $\ast$ The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (goerg2013forecastable). A larger value indicates better predictability.
 D.2 Baseline Details
-We employ five state-of-the-art Transformer-based models, such as iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), WPMixer (Murad et al., 2025), and MultiPatchFormer (Naghashi et al., 2025), as well as three well-acknowledged non-Transformer-based models, Dlinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023).
-We utilized the official repositories of the Time-Series-Library (Wang et al., 2024b) directly for reproduction, and adopted the same comparison settings.
+We employ four state-of-the-art Transformer-based models, such as iTransformer (liu2024itransformer), PatchTST (nie2023time), FEDformer (zhou2022fedformer), and MultiPatchFormer (naghashi2025multipatchformer), two well-acknowledged Linear models, Dlinear (zeng2023dlinear), Nlinear, one prominent RNN-based model SegRNN (lin2023segrnn) and WPMixer (murad2024wp).
+We comprehensively evaluate the HTMformer framework using four mainstream attention mechanisms Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention).
+We utilized the official repositories of the Time-Series-Library (wang2024tssurvey) directly for reproduction, and adopted the same comparison settings.
 The Time-Series-Library consolidates and integrates implementations of prior time-series forecasting models, making it convenient for us to conduct experiments.
 Among these models, MultiPatchFormer is a highly advanced architecture and yields the best overall forecasting accuracy.
-Both iTransformer and our model utilize input inversion to enable the attention mechanism to effectively capture multivariate correlations.
-WPMixer efficiently captures both the frequency-domain and time-domain information in time series data, enabling deep temporal feature extraction.
-FEDformer achieves high forecasting accuracy and computational efficiency through an innovative architecture that integrates seasonal-trend decomposition with frequency-domain enhancement.
-Meanwhile, DLinear is a state-of-the-art linear model with a non-Transformer architecture.
-SegRNN has demonstrated outstanding performance on multiple benchmark time series forecasting datasets by employing segment-wise iterative processing and parallelized multi-step prediction.
-TimeMixer is a linear time series forecasting model that leverages a multi-scale decomposable mixing mechanism, demonstrating robust performance on high-dimensional datasets.
+iTransformer and HTMformers share the same architecture.
+HTMformers replaces the embedding layer of iTransformer.
 Note that SegRNN is incompatible with short-term forecasting, so no comparison is conducted.
 D.3 Experiment Details
-All experiments were implemented in PyTorch (Paszke et al., 2019) and conducted on a dedicated NVIDIA Quadro RTX 6000 GPU.
-We trained the model for 10 epochs using the Adam optimizer (Kingma & Ba, 2015), with an early stopping mechanism configured at a patience value of 3.
-The key parameter configurations are listed in Table 7 (WPMixer is configured with d_model = 16 and d_ff = 32).
-We verified that, for these baselines, all other hyperparameters were taken from their official repositories within the Time-Series-Library (Wang et al., 2024b), which consolidates implementations of prior time-series forecasting models.
+All experiments were implemented in PyTorch (paszke2019pytorch) and conducted on a dedicated NVIDIA Quadro RTX 6000 GPU.
+We trained the model for 10 epochs using the Adam optimizer (kingma2015adam), with an early stopping mechanism configured at a patience value of 3.
+The key parameter configurations are listed in Table 6.
+We verified that, for these baselines, all other hyperparameters were taken from their official repositories within the Time-Series-Library (wang2024tssurvey), which consolidates implementations of prior time-series forecasting models.
 This ensures consistency with the established fair comparison protocol.
@@
 E.1 Performance Promotion with HTME
-To assess the effectiveness and scalability of the HTME for enhancing the Transformer’s capacity to learn temporal sequence representations, we integrated HTME into the Transformer and its variants: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022).
+To assess the effectiveness and scalability of the HTMformers for enhancing the Attention’s capacity to learn temporal sequence representations, we integrated various attention variants: Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention).
+We apply the HTME strategy to models with other architectures, including the Transformer-based iTransformer, the linear-based DLinear, NLinear and the RNN-based SegRNN.
 The length of historical data sequences was set to 96 time steps, with the forecast horizon varying among 96, 192, 336, and 720 time steps.
 The full predictive outcomes are detailed in Table 7.
-On three benchmark datasets (ECL, Traffic, and Weather), the HTME versions yield significant improvements over all four models across all proposed prediction horizons, thereby confirming their effectiveness and scalability.
-We further integrated the multivariate feature extraction module from the HTME extractor into iTransformer to assess the scalability capability of this module, with the detailed results presented in the ablation study.
-This component was seamlessly incorporated into the iTransformer’s embedding layer, leading to a substantial improvement in its overall performance.
+On three benchmark datasets (ECL, Traffic, and Weather), the HTME versions yield significant improvements over all four attentions across all proposed prediction horizons.
+Across three mainstream forecasting architectures, the HTME strategy consistently improves performance by margins ranging from 1.7% to 4.6%.
+These results highlight the effectiveness and robustness of the HTME strategy.
 Table 7: Full results of original Transformers and Transformers with HTME.
-Models Transformer Reformer Informer Flowformer Flashformer
-Dataset Type Step MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Electricity Original 96 0.264 0.363 0.304 0.391 0.330 0.415 0.272 0.371 0.263 0.361
-192 0.275 0.375 0.350 0.426 0.356 0.438 0.269 0.371 0.270 0.369
-336 0.262 0.362 0.354 0.429 0.376 0.454 0.264 0.364 0.263 0.360
-720 0.286 0.373 0.361 0.422 0.454 0.464 0.284 0.374 0.279 0.368
-Avg 0.271 0.363 0.342 0.417 0.379 0.442 0.272 0.370 0.268 0.364
-+HTME 96 0.157 0.247 0.169 0.266 0.170 0.266 0.170 0.266 0.163 0.262
-192 0.171 0.260 0.190 0.283 0.187 0.279 0.188 0.280 0.183 0.277
-336 0.192 0.277 0.200 0.293 0.207 0.296 0.202 0.294 0.199 0.292
-720 0.220 0.304 0.248 0.329 0.247 0.329 0.235 0.319 0.229 0.316
-Avg 0.185 0.272 0.201 0.292 0.202 0.292 0.198 0.289 0.194 0.287
-Weather Original 96 0.372 0.417 0.311 0.355 0.420 0.452 0.283 0.364 0.326 0.395
-192 0.632 0.580 0.453 0.464 0.423 0.455 0.548 0.541 0.587 0.544
-336 0.699 0.606 0.476 0.474 0.588 0.539 0.614 0.548 0.714 0.614
-720 0.903 0.688 0.579 0.541 1.146 0.784 1.079 0.783 0.906 0.684
-Avg 0.651 0.572 0.454 0.458 0.644 0.557 0.631 0.559 0.633 0.559
-+HTME 96 0.163 0.209 0.164 0.209 0.170 0.213 0.167 0.212 0.168 0.215
-192 0.222 0.261 0.214 0.255 0.220 0.260 0.216 0.255 0.218 0.258
-336 0.272 0.290 0.270 0.288 0.274 0.298 0.279 0.302 0.275 0.298
-720 0.359 0.352 0.351 0.348 0.354 0.351 0.354 0.350 0.352 0.349
-Avg 0.254 0.277 0.249 0.275 0.254 0.280 0.254 0.279 0.253 0.380
-Traffic Original 96 0.658 0.364 0.715 0.399 0.731 0.413 0.646 0.359 0.637 0.353
-192 0.655 0.359 0.705 0.390 0.761 0.431 0.646 0.355 0.659 0.364
-336 0.657 0.361 0.703 0.386 0.840 0.474 0.657 0.358 0.653 0.357
-720 0.697 0.380 0.703 0.383 0.969 0.545 0.671 0.364 0.697 0.375
-Avg 0.666 0.366 0.706 0.389 0.825 0.465 0.655 0.359 0.661 0.362
-+HTME 96 0.439 0.300 0.445 0.302 0.466 0.314 0.451 0.309 0.439 0.300
-192 0.450 0.302 0.463 0.308 0.468 0.312 0.466 0.310 0.460 0.310
-336 0.462 0.310 0.478 0.313 0.488 0.322 0.484 0.320 0.480 0.321
-720 0.517 0.339 0.514 0.334 0.530 0.347 0.524 0.345 0.516 0.342
-Avg 0.467 0.312 0.475 0.314 0.488 0.323 0.481 0.321 0.474 0.319
+Models Transformer Reformer Informer Flowformer Flashformer iTransformer DLinear Nlinear SegRNN
+Dataset Type Step MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+Electricity Original 96 0.264 0.363 0.304 0.391 0.330 0.415 0.272 0.371 0.263 0.361 0.163 0.252 0.210 0.301 0.199 0.277 0.168 0.264
+192 0.275 0.375 0.350 0.426 0.356 0.438 0.269 0.371 0.270 0.369 0.175 0.263 0.210 0.304 0.199 0.280 0.183 0.279
+336 0.262 0.362 0.354 0.429 0.376 0.454 0.264 0.364 0.263 0.360 0.282 0.299 0.223 0.319 0.214 0.295 0.192 0.289
+720 0.286 0.373 0.361 0.422 0.454 0.464 0.284 0.374 0.279 0.368 0.357 0.348 0.257 0.349 0.256 0.328 0.231 0.321
+Avg 0.271 0.363 0.342 0.417 0.379 0.442 0.272 0.370 0.268 0.364 0.244 0.290 0.225 0.318 0.217 0.295 0.194 0.288
++HTME 96 0.157 0.247 0.169 0.266 0.170 0.266 0.170 0.266 0.163 0.262 0.157 0.247 0.204 0.299 0.189 0.273 0.164 0.258
+192 0.171 0.260 0.190 0.283 0.187 0.279 0.188 0.280 0.183 0.277 0.171 0.260 0.211 0.307 0.192 0.273 0.175 0.269
+336 0.192 0.277 0.200 0.293 0.207 0.296 0.202 0.294 0.199 0.292 0.192 0.277 0.222 0.318 0.206 0.290 0.192 0.285
+720 0.220 0.304 0.248 0.329 0.247 0.329 0.235 0.319 0.229 0.316 0.220 0.304 0.24 0.331 0.246 0.329 0.230 0.320
+Avg 0.185 0.272 0.201 0.292 0.202 0.292 0.198 0.289 0.194 0.287 0.185 0.272 0.219 0.314 0.208 0.291 0.190 0.283
+Weather Original 96 0.372 0.417 0.311 0.355 0.420 0.452 0.283 0.364 0.326 0.395 0.176 0.216 0.193 0.233 0.196 0.237 0.168 0.23
+192 0.632 0.580 0.453 0.464 0.423 0.455 0.548 0.541 0.587 0.544 0.223 0.255 0.236 0.268 0.241 0.279 0.215 0.277
+336 0.699 0.606 0.476 0.474 0.588 0.539 0.614 0.548 0.714 0.614 0.280 0.298 0.288 0.304 0.293 0.317 0.274 0.319
+720 0.903 0.688 0.579 0.541 1.146 0.784 1.079 0.783 0.906 0.684 0.353 0.346 0.359 0.349 0.366 0.361 0.359 0.378
+Avg 0.651 0.572 0.454 0.458 0.644 0.557 0.631 0.559 0.633 0.559 0.259 0.279 0.269 0.289 0.274 0.299 0.254 0.301
++HTME 96 0.163 0.209 0.164 0.209 0.170 0.213 0.167 0.212 0.168 0.215 0.163 0.209 0.177 0.223 0.168 0.230 0.161 0.226
+192 0.222 0.261 0.214 0.255 0.220 0.260 0.216 0.255 0.218 0.258 0.222 0.261 0.227 0.266 0.224 0.275 0.211 0.273
+336 0.272 0.290 0.270 0.288 0.274 0.298 0.279 0.302 0.275 0.298 0.272 0.290 0.281 0.303 0.295 0.313 0.269 0.315
+720 0.359 0.352 0.351 0.348 0.354 0.351 0.354 0.350 0.352 0.349 0.359 0.352 0.358 0.352 0.359 0.343 0.355 0.371
+Avg 0.254 0.277 0.249 0.275 0.254 0.280 0.254 0.279 0.253 0.380 0.254 0.277 0.261 0.286 0.261 0.290 0.249 0.296
+Traffic Original 96 0.658 0.364 0.715 0.399 0.731 0.413 0.646 0.359 0.637 0.353 0.442 0.302 0.696 0.428 0.654 0.396 0.723 0.376
+192 0.655 0.359 0.705 0.390 0.761 0.431 0.646 0.355 0.659 0.364 0.459 0.308 0.646 0.407 0.607 0.374 0.737 0.385
+336 0.657 0.361 0.703 0.386 0.840 0.474 0.657 0.358 0.653 0.357 0.479 0.319 0.653 0.409 0.615 0.376 0.766 0.405
+720 0.697 0.380 0.703 0.383 0.969 0.545 0.671 0.364 0.697 0.375 0.516 0.342 0.694 0.428 0.654 0.417 0.818 0.423
+Avg 0.666 0.366 0.706 0.389 0.825 0.465 0.655 0.359 0.661 0.362 0.474 0.317 0.672 0.418 0.632 0.391 0.761 0.397
++HTME 96 0.439 0.300 0.445 0.302 0.466 0.314 0.451 0.309 0.439 0.300 0.439 0.300 0.626 0.393 0.624 0.394 0.696 0.366
+192 0.450 0.302 0.463 0.308 0.468 0.312 0.466 0.310 0.460 0.310 0.450 0.302 0.583 0.369 0.588 0.372 0.717 0.382
+336 0.462 0.310 0.478 0.313 0.488 0.322 0.484 0.320 0.480 0.321 0.462 0.310 0.591 0.373 0.597 0.376 0.754 0.401
+720 0.517 0.339 0.514 0.334 0.530 0.347 0.524 0.345 0.516 0.342 0.517 0.339 0.636 0.395 0.635 0.395 0.803 0.419
+Avg 0.467 0.312 0.475 0.314 0.488 0.323 0.481 0.321 0.474 0.319 0.467 0.312 0.609 0.383 0.611 0.384 0.750 0.392
 E.2 Long-term Forecasting Results
@@
 SegRNN
+nie2023time
+zeng2023dlinear
+zhou2022fedformer
+naghashi2025multipatchformer
+murad2024wp
+wang2023timemixer
+lin2023segrnn
 Metric
@@
 TimeMixer
+nie2023time
+zeng2023dlinear
+zhou2022fedformer
+naghashi2025multipatchformer
+murad2024wp
+wang2023timemixer
 Metric
@@
 Appendix F Additional Ablation Studies
-To investigate the functional rationale of HTMformer components, we conduct detailed ablation experiments involving both component replacement and component removal.
+To investigate the functional rationale of HTME components, We examined three different variants of iTransformer, DLinear, and SegRNN, as representatives of Transformer-based, Linear-based, and RNN-based models, respectively.
+1) Productor, which integrates the multivariate feature extraction module of the HTME extractor into the embedding layer of iTransformer.
+2) ProductorV1, which utilizes only the multivariate feature extraction module of the HTME extractor.
+We also examined another variant 3) ProductorV2, which utilizes the multivariate feature extraction module of the HTME extractor to supplement the embedding representations.
+We introduce two Additional variants: HTMformer, builds upon iTransformer by incorporating the complete HTME extractor; HTMformerV1, replaces the original embedding layer in iTransformer with the temporal feature extractor from HTME.
 The detailed results are reported in Table 12 of the main text.
-As all experiments were conducted under identical configurations, the HTMformer variants are directly comparable to other models.
+As all experiments were conducted under identical configurations, the HTMPredictors are directly comparable to other models.
 Effect of the temporal feature extraction module.
@@
 Observations indicate that HTMformerV1 demonstrates improvements in 63/80 cases.
-By comparing HTMformerV1 and HTMformerV2, numbers prefixed with an underline indicate that HTMformerV1 outperforms HTMformerV2.
-Notably, HTMformerV1 performs better in 57/80 cases.
-Significantly, on Solar and PEMS datasets lacking timestamps, HTMformerV2 often exhibits superior performance, verifying our first hypothesis: 1) The majority of time series features are stored in the temporal dimension, and the depth of temporal feature mining substantially influences prediction accuracy.
+Effect of the temporal feature.
+By comparing Predictor and PredictorV2, numbers prefixed with an underline indicate that Predictor outperforms PredictorV2.
+Notably, Predictor performs better in most cases.
+The majority of time series features are stored in the temporal dimension, and the depth of temporal feature mining substantially influences prediction accuracy.
+This indicates that the temporal feature extraction module of HTME is more effective.
 Effect of the multivariate features.
-Using iTransformer and HTMformerV1 as baselines, we compare them against iTransformerV3 and HTMformer versions.
+Using Predictor as baselines, we compare them against PredictorV1.
 Red highlighted numbers indicate cases where the multivariate-integrated versions outperform their original counterparts.
-Experimental results show that iTransformerV3 achieves superior performance in 61/80 cases compared to iTransformer, while HTMformer outperforms its baseline in 68/80 cases.
-The gains are particularly evident on the PEMS datasets, which are derived from public transportation network data characterized by strong correlations among sensor nodes.
-This finding corroborates our second hypothesis: 2) The spatial dimension (multivariate dimension) also contains abundant features, and comprehensive multivariate feature mining can further reduce errors.
+Experimental results show that PredictorV1 achieves superior performance most cases.
+We also find that PredictorV2 outperforms Predictor in some cases.
+The spatial dimension (multivariate dimension) also contains abundant features, and comprehensive multivariate feature mining can further reduce errors.
 Effect of the hybrid strategy.
-Among all models and their variants, HTMformer excels in 68/80 cases.
-This demonstrates the effectiveness of the combination.
+Among all models and their variants, PredictorV1 achieves superior performance.
+This demonstrates the effectiveness of the HTME strategy.
 In most cases, temporal features are the primary characteristics of time series data, but the correlations between multiple variables cannot be overlooked.
-It is of great significance for improving the accuracy of Transformer-based predictors.
-Different datasets vary in the depth of temporal and multivariate features.
-Therefore, we use a learnable parameter to control the proportion of the two modules, allowing the fusion of multivariate features to still reduce errors to an extent.
-In conclusion, the strategy of fusing temporal features and multivariate features is superior to using either of them alone.
-This strategy can enhance the accuracy and scalability of Transformer-based predictors.
+The HTME strategy has successfully suppressed noise and extracted effective patterns from the raw sequence.
+In conclusion, the strategy of fusing temporal features and multivariate features can enhance the accuracy and scalability of Transformer-based predictors.
 Table 12: We compare different variants under different prediction lengths on multiple datasets (MSE and MAE). The input sequence length is set to 96.
@@
 HTMformerV1
-HTMformerV2
+iTransV2
 iTransformer
-iTransV3
-MPFormer
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ECL 96 0.157 0.247 0.153 0.246 0.255 0.34 0.163 0.252 0.165 0.262 0.159 0.249
-192 0.171 0.260 0.169 0.260 0.264 0.349 0.175 0.263 0.184 0.278 0.171 0.260
-336 0.192 0.277 0.187 0.277 0.273 0.357 0.282 0.299 0.199 0.294 0.188 0.277
-720 0.220 0.304 0.227 0.311 0.31 0.384 0.357 0.348 0.231 0.318 0.228 0.309
-AVG 0.185 0.272 0.184 0.273 0.275 0.357 0.244 0.29 0.194 0.288 0.186 0.273
-Weather 96 0.164 0.209 0.169 0.212 0.175 0.229 0.176 0.216 0.167 0.167 0.168 0.209
-192 0.222 0.261 0.218 0.256 0.223 0.264 0.223 0.255 0.217 0.256 0.213 0.250
-336 0.272 0.290 0.275 0.298 0.287 0.31 0.28 0.298 0.272 0.297 0.273 0.293
-720 0.359 0.352 0.353 0.349 0.364 0.358 0.357 0.349 0.353 0.349 0.354 0.347
-AVG 0.254 0.277 0.253 0.278 0.262 0.29 0.259 0.279 0.252 0.267 0.252 0.274
-Traffic 96 0.439 0.300 0.447 0.303 0.648 0.405 0.442 0.302 0.440 0.309 0.433 0.290
-192 0.450 0.302 0.460 0.307 0.737 0.438 0.459 0.308 0.457 0.305 0.450 0.296
-336 0.462 0.310 0.482 0.319 0.745 0.435 0.479 0.319 0.478 0.321 0.467 0.306
-720 0.517 0.339 0.522 0.343 0.863 0.484 0.516 0.342 0.520 0.349 0.500 0.326
-AVG 0.467 0.312 0.477 0.318 0.748 0.44 0.474 0.317 0.473 0.321 0.462 0.304
-ETTh2 96 0.300 0.348 0.298 0.348 0.375 0.399 0.295 0.344 0.304 0.352 0.299 0.350
-192 0.389 0.402 0.387 0.402 0.45 0.435 0.375 0.398 0.391 0.406 0.384 0.401
-336 0.409 0.408 0.416 0.427 0.481 0.469 0.436 0.435 0.419 0.428 0.411 0.428
-720 0.421 0.439 0.426 0.443 0.471 0.474 0.429 0.447 0.422 0.438 0.430 0.446
-AVG 0.379 0.399 0.381 0.405 0.444 0.444 0.383 0.406 0.384 0.406 0.381 0.406
-Solar 96 0.199 0.239 0.209 0.252 0.244 0.284 0.213 0.253 0.207 0.249 0.203 0.240
-192 0.233 0.263 0.238 0.274 0.277 0.306 0.242 0.274 0.240 0.279 0.237 0.268
-336 0.252 0.279 0.261 0.292 0.323 0.326 0.262 0.29 0.256 0.290 0.249 0.280
-720 0.259 0.280 0.265 0.295 0.311 0.329 0.27 0.296 0.275 0.303 0.254 0.283
-AVG 0.235 0.265 0.243 0.278 0.288 0.311 0.246 0.278 0.244 0.280 0.235 0.267
-PMES03 96 0.250 0.343 0.304 0.393 0.302 0.363 0.337 0.413 0.240 0.340 0.302 0.395
-192 0.297 0.378 0.353 0.424 0.373 0.422 0.374 0.436 0.300 0.384 0.349 0.428
-336 0.274 0.356 0.318 0.388 0.32 0.388 0.332 0.395 0.286 0.365 0.313 0.388
-720 0.336 0.401 0.387 0.438 0.409 0.44 0.397 0.44 0.352 0.414 0.383 0.441
-AVG 0.289 0.369 0.340 0.410 0.351 0.403 0.36 0.421 0.294 0.375 0.336 0.413
-PMES04 96 0.226 0.335 0.364 0.433 0.236 0.348 0.387 0.446 0.235 0.344 0.356 0.428
-192 0.286 0.383 0.406 0.406 0.326 0.413 0.429 0.473 0.307 0.395 0.398 0.459
-336 0.286 0.372 0.359 0.417 0.312 0.398 0.369 0.426 0.297 0.378 0.348 0.413
-720 0.338 0.414 0.436 0.471 0.387 0.451 0.44 0.476 0.362 0.428 0.423 0.467
-AVG 0.284 0.376 0.391 0.431 0.315 0.402 0.406 0.455 0.300 0.386 0.381 0.441
-PMES08 96 0.365 0.409 0.469 0.461 0.401 0.416 0.477 0.47 0.338 0.388 0.406 0.437
-192 0.513 0.462 0.612 0.507 0.567 0.481 0.627 0.518 0.487 0.444 0.539 0.479
-336 0.530 0.435 0.606 0.471 0.567 0.452 0.599 0.472 0.516 0.425 0.538 0.444
-720 0.624 0.487 0.692 0.531 0.697 0.512 0.691 0.532 0.600 0.479 0.603 0.489
-AVG 0.508 0.448 0.594 0.492 0.558 0.465 0.598 0.498 0.485 0.434 0.521 0.462
+iTransV1
+DLinearV2
+DLinear
+DLinearV1
+SegRNNV2
+SegRNN
+SegRNNV1
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ECL 96 0.157 0.247 0.153 0.246 0.255 0.34 0.163 0.252 0.165 0.262 0.499 0.533 0.210 0.301 0.204 0.299 0.390 0.454 0.168 0.264 0.164 0.258
+192 0.171 0.260 0.169 0.260 0.264 0.349 0.175 0.263 0.184 0.278 0.505 0.535 0.210 0.304 0.211 0.307 0.407 0.464 0.183 0.279 0.175 0.269
+336 0.192 0.277 0.187 0.277 0.273 0.357 0.282 0.299 0.199 0.294 0.504 0.531 0.223 0.319 0.222 0.318 0.415 0.470 0.192 0.289 0.192 0.285
+720 0.220 0.304 0.227 0.311 0.31 0.384 0.357 0.348 0.231 0.318 0.491 0.526 0.257 0.349 0.240 0.331 0.452 0.491 0.231 0.321 0.230 0.320
+AVG 0.185 0.272 0.184 0.273 0.275 0.357 0.244 0.29 0.194 0.288 0.500 0.531 0.225 0.318 0.219 0.314 0.416 0.470 0.194 0.288 0.190 0.283
+Weather 96 0.164 0.209 0.169 0.212 0.175 0.229 0.176 0.216 0.167 0.167 0.18 0.272 0.193 0.233 0.177 0.223 0.165 0.231 0.168 0.230 0.161 0.226
+192 0.222 0.261 0.218 0.256 0.223 0.264 0.223 0.255 0.217 0.256 0.232 0.320 0.236 0.268 0.227 0.266 0.221 0.285 0.215 0.277 0.211 0.273
+336 0.272 0.290 0.275 0.298 0.287 0.31 0.28 0.298 0.272 0.297 0.294 0.365 0.288 0.304 0.281 0.303 0.281 0.330 0.274 0.319 0.269 0.315
+720 0.359 0.352 0.353 0.349 0.364 0.358 0.357 0.349 0.353 0.349 0.391 0.433 0.359 0.349 0.358 0.352 0.364 0.385 0.359 0.378 0.355 0.371
+AVG 0.254 0.277 0.253 0.278 0.262 0.29 0.259 0.279 0.252 0.267 0.274 0.348 0.269 0.289 0.261 0.286 0.258 0.308 0.254 0.301 0.249 0.296
+Traffic 96 0.439 0.300 0.447 0.303 0.648 0.405 0.442 0.302 0.440 0.309 0.955 0.544 0.696 0.428 0.626 0.393 1.224 0.583 0.723 0.376 0.696 0.366
+192 0.450 0.302 0.460 0.307 0.737 0.438 0.459 0.308 0.457 0.305 0.970 0.543 0.646 0.407 0.583 0.369 1.201 0.568 0.737 0.385 0.717 0.382
+336 0.462 0.310 0.482 0.319 0.745 0.435 0.479 0.319 0.478 0.321 0.982 0.550 0.653 0.409 0.591 0.373 1.285 0.595 0.766 0.405 0.754 0.401
+720 0.517 0.339 0.522 0.343 0.863 0.484 0.516 0.342 0.520 0.349 0.997 0.548 0.694 0.428 0.636 0.395 1.343 0.614 0.818 0.423 0.803 0.419
+AVG 0.467 0.312 0.477 0.318 0.748 0.44 0.474 0.317 0.473 0.321 0.976 0.546 0.672 0.418 0.609 0.383 1.263 0.590 0.761 0.397 0.750 0.392
+ETTh2 96 0.300 0.348 0.298 0.348 0.375 0.399 0.295 0.344 0.304 0.352 0.357 0.409 0.341 0.395 0.343 0.399 0.339 0.389 0.295 0.355 0.292 0.345
+192 0.389 0.402 0.387 0.402 0.45 0.435 0.375 0.398 0.391 0.406 0.497 0.491 0.481 0.479 0.480 0.477 0.417 0.431 0.375 0.397 0.371 0.396
+336 0.409 0.408 0.416 0.427 0.481 0.469 0.436 0.435 0.419 0.428 0.610 0.553 0.592 0.542 0.597 0.540 0.476 0.476 0.414 0.436 0.416 0.430
+720 0.421 0.439 0.426 0.443 0.471 0.474 0.429 0.447 0.422 0.438 0.816 0.658 0.840 0.661 0.820 0.652 0.474 0.489 0.443 0.469 0.436 0.462
+AVG 0.379 0.399 0.381 0.405 0.444 0.444 0.383 0.406 0.384 0.406 0.570 0.528 0.563 0.519 0.560 0.517 0.427 0.446 0.382 0.414 0.379 0.408
+Solar 96 0.199 0.239 0.209 0.252 0.244 0.284 0.213 0.253 0.207 0.249 0.231 0.326 0.289 0.377 0.225 0.317 0.353 0.416 0.232 0.294 0.203 0.291
+192 0.233 0.263 0.238 0.274 0.277 0.306 0.242 0.274 0.240 0.279 0.262 0.348 0.319 0.397 0.260 0.342 0.404 0.448 0.253 0.306 0.230 0.303
+336 0.252 0.279 0.261 0.292 0.323 0.326 0.262 0.29 0.256 0.290 0.302 0.373 0.352 0.415 0.296 0.367 0.418 0.455 0.264 0.310 0.257 0.321
+720 0.259 0.280 0.265 0.295 0.311 0.329 0.27 0.296 0.275 0.303 0.313 0.378 0.356 0.412 0.309 0.373 0.482 0.516 0.260 0.306 0.281 0.349
+AVG 0.235 0.265 0.243 0.278 0.288 0.311 0.246 0.278 0.244 0.280 0.277 0.356 0.329 0.400 0.273 0.350 0.414 0.459 0.252 0.304 0.243 0.316
+PMES03 96 0.250 0.343 0.304 0.393 0.302 0.363 0.337 0.413 0.240 0.340 0.315 0.418 0.458 0.517 0.313 0.416 0.389 0.446 0.381 0.438 0.287 0.374
+192 0.297 0.378 0.353 0.424 0.373 0.422 0.374 0.436 0.300 0.384 0.410 0.477 0.477 0.527 0.410 0.477 0.620 0.602 0.447 0.478 0.368 0.433
+336 0.274 0.356 0.318 0.388 0.32 0.388 0.332 0.395 0.286 0.365 0.337 0.422 0.396 0.456 0.344 0.420 0.729 0.641 0.380 0.432 0.349 0.418
+720 0.336 0.401 0.387 0.438 0.409 0.44 0.397 0.44 0.352 0.414 0.391 0.469 0.439 0.492 0.396 0.462 0.774 0.683 0.447 0.478 0.406 0.472
+AVG 0.289 0.369 0.340 0.410 0.351 0.403 0.36 0.421 0.294 0.375 0.363 0.447 0.442 0.498 0.366 0.444 0.628 0.593 0.413 0.456 0.353 0.424
+PMES04 96 0.226 0.335 0.364 0.433 0.236 0.348 0.387 0.446 0.235 0.344 0.315 0.404 0.452 0.504 0.312 0.401 0.370 0.457 0.415 0.468 0.291 0.398
+192 0.286 0.383 0.406 0.406 0.326 0.413 0.429 0.473 0.307 0.395 0.377 0.451 0.477 0.527 0.372 0.451 0.723 0.665 0.496 0.510 0.463 0.507
+336 0.286 0.372 0.359 0.417 0.312 0.398 0.369 0.426 0.297 0.378 0.330 0.411 0.396 0.456 0.330 0.410 0.789 0.694 0.420 0.461 0.453 0.508
+720 0.338 0.414 0.436 0.471 0.387 0.451 0.44 0.476 0.362 0.428 0.401 0.460 0.439 0.492 0.400 0.460 0.850 0.729 0.479 0.502 0.714 0.677
+AVG 0.284 0.376 0.391 0.431 0.315 0.402 0.406 0.455 0.300 0.386 0.356 0.432 0.441 0.494 0.354 0.431 0.683 0.636 0.452 0.485 0.480 0.522
+PMES08 96 0.365 0.409 0.469 0.461 0.401 0.416 0.477 0.47 0.338 0.388 0.449 0.446 0.672 0.564 0.454 0.451 0.429 0.477 0.516 0.487 0.364 0.425
+192 0.513 0.462 0.612 0.507 0.567 0.481 0.627 0.518 0.487 0.444 0.599 0.536 0.726 0.580 0.597 0.535 0.811 0.674 0.646 0.528 0.568 0.523
+336 0.530 0.435 0.606 0.471 0.567 0.452 0.599 0.472 0.516 0.425 0.569 0.476 0.650 0.518 0.564 0.469 0.949 0.712 0.617 0.490 0.621 0.511
+720 0.624 0.487 0.692 0.531 0.697 0.512 0.691 0.532 0.600 0.479 0.633 0.531 0.713 0.562 0.632 0.521 1.034 0.743 0.707 0.541 0.874 0.652
+AVG 0.508 0.448 0.594 0.492 0.558 0.465 0.598 0.498 0.485 0.434 0.563 0.497 0.690 0.556 0.562 0.494 0.806 0.652 0.621 0.511 0.607 0.528
 Appendix G Details of Model Efficiency
-For a comprehensive efficiency comparison, we evaluate HTMformer against three highly competitive baselines, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023), across eight datasets.
-The evaluation considers three key metrics: training time, GPU memory footprint, and total parameter count, with the input sequence length fixed at 96 and the prediction length at 192.
-Moreover, the experimental hardware conditions and parameter configurations follow those described in Appendix B, and the best results are marked in red.
+For a comprehensive efficiency comparison, we evaluate HTMformer against three highly competitive baselines, including MultiPatchFormer (naghashi2025multipatchformer), WPMixer (murad2024wp), and PatchTST (nie2023time), across eight datasets.
+The evaluation considers three key metrics: training time, GPU memory footprint, with the input sequence length fixed at 96 and the prediction length at 192.
+Moreover, the experimental hardware conditions and parameter configurations follow those described in Appendix D, and the best results are marked in red.
 The full results are shown in Table 13 (Left) and Table 13 (Right).
-The training time and GPU usage serve as an intuitive indicator of total parameter count, so detailed reporting is omitted.
 Across all datasets, HTMformer achieves the best training time and GPU memory footprint.
@@
 Moreover, this advantage becomes increasingly pronounced as dimensionality grows.
-This superiority stems from our lightweight design philosophy, in which HTME avoids overemphasizing temporal feature extraction and HTMformer adopts only a vanilla Transformer encoder combined with the inverted input strategy.
+HTME is a lightweight module that enables simple models focused on temporal feature modeling to incorporate the influence of multivariate correlations, thereby achieving performance comparable to that of complex models.
 […]
 Datasets HTMFormer MPFormer WPMixer PatchTST
-Models Ours 2025 2025 2023
+Models Ours naghashi2025multipatchformer murad2024wp nie2023time
 ECL 0.1004 0.3319 0.4143 0.5088
@@
 Datasets HTMFormer MPFormer WPMixer PatchTST
-Models Ours 2025 2025 2023
+Models Ours naghashi2025multipatchformer murad2024wp nie2023time
 ECL 1623.47 7471.94 13784.07 7195.99
@@
 AVG 1613.82 6483.46 10125.27 6152.02
-Appendix H Hyperparameter Sensitivity
+Appendix H Analysis of HTME
+The vast majority of features in time series are contained within the temporal dimension, and the depth of exploration in this dimension significantly influences prediction accuracy (hyndman2021forecasting).
+Therefore, existing models can achieve superior performance even by focusing solely on the temporal dimension.
+Moreover, thanks to their simple model architectures, they often have a low computational overhead.
+The multivariate dimension also contains a large number of features (Pravilovic2013SpatialForecasting), and fully exploiting multivariate features can further boost performance.
+For time series data, the vast majority of features are latent within the temporal dimension, making the deep extraction of temporal features the key to effective modeling.
+Overemphasizing multivariate features—particularly strategies that explicitly model cross-variable correlations within the backbone network—introduces significant computational overhead without a corresponding performance gain.
+Most importantly, introducing additional complex modules into the backbone network to model multivariate correlations forces temporal and multivariate features to share the same feature space.
+Consequently, while the model is capturing cross-variable relationships, the representation of temporal features is inevitably compromised.
+We visualize the temporal and multivariate feature spaces of HTME, as well as the weight parameter alpha.
+The value of alpha is approximately 0.65, indicating that subsequent modules are more inclined to select temporal features.
+Additionally, temporal feature weights are considerably higher than spatial feature weights.
+This emphasizes the primary role of temporal information.
+We also observe that the temporal features exhibit a strip-like distribution, showing significant variation along the temporal dimension but minimal differences in the multivariate dimension.
+This suggests that the temporal features possess a high degree of consistency across the multivariate dimension, which in turn underscores their importance.
+The multivariate features exhibit a point-like and uniform distribution.
+For the most part, they take on small values, which ensures that they do not disrupt the temporal feature space.
+A few points within the multivariate features exhibit large values.
+These points serve to emphasize their corresponding temporal features, thereby creating a more information-rich temporal feature space.
+Consequently, subsequent modules can utilize this enhanced feature space more efficiently, improving the model’s capability to capture complex temporal dynamics.
+Appendix I Hyperparameter Sensitivity
 We assess the sensitivity of HTMformer to variations in key hyperparameters, including the learning rate $lr$, the number of Transformer blocks $L$, the FCN dimension $K$, and the hidden dimension $D$ of the variate tokens.
@@
 Figure 7: Hyperparameter sensitivity with respect to lrlr, $L$, $K$, and $D$ . All results are obtained using a lookback window of $T=96$ time steps and a forecast horizon of $S=96$ time steps.
-Appendix I Visualization of Prediction Results
-We visualize HTMformer’s ability to predict trends across various time series datasets, including ECL, Traffic, Solar-Energy, Weather, and PEMS (PEMS03, PEMS04 and PEMS08), as shown in Figures˜7, 8, 9, 10, 11, 12, 13 and 14.
+Appendix J Visualization of Prediction Results
+We visualize HTMformer’s ability to predict trends across various time series datasets, including ECL, Traffic, Solar-Energy, Weather, and PEMS (PEMS03, PEMS04 and PEMS08), as shown in Figures˜8, 9, 10, 11, 12, 13, 14 and 15.
 Each example uses a 96-step input to generate 96-step predictions.
@@
 Prediction showcases are listed in Figure 16 and Figure 17.
-Appendix J Discussion on Transformer-based Predictors
-Recent studies (Lu et al., 2023; Li et al., 2023b) have found that Channel Dependence ideally gains from higher capacity, while Channel Independence can significantly enhance performance due to sample scarcity.
-In time series forecasting, an increasing number of models employ Channel Independence (Nie et al., 2023), as most current forecasting benchmarks are not sufficiently large.
-These models treat time series variates independently and utilize a shared backbone, achieving good performance.
-However, Channel Independence inevitably overlooks the multivariate correlations that exist in most datasets, resulting in the loss of some information.
-We believe that making variates independent is desirable, but multivariate correlations should not be ignored.
-However, previous improvements to models have mainly focused on extracting features in the temporal dimension.
-Excessive extraction of temporal features does not further improve prediction accuracy, but rather significantly increases computational overhead.
-Notably, as demonstrated in both the introduction and ablation studies of this paper.
-The main reason why traditional Transformer-based predictors encounter performance bottlenecks is that the embedding model extracts only limited effective information from the sequence, which in turn affects the performance of subsequent models.
-Due to the inherent limitations of Transformer-based predictors, simply enlarging the input sequence length cannot further enhance accuracy.
-Constructing an embedding model that deeply extracts sequence information can effectively address this problem.
-Therefore, we propose a novel strategy, specifically extracting multivariate features in the embedding layer and using them to complement temporal features.
-This approach incorporates multivariate correlations, thereby further extracting effective information from the sequence.
-Moreover, it does not affect the subsequent attention mechanism to utilize Channel Independence.
-We find that extracting local patterns in time series through patching and convolution, together with capturing long-term dependencies using linear layers, is sufficient for effectively modeling temporal features.
-For Transformer-based predictors, this is sufficient to extract temporal representations, and it is unnecessary to employ attention mechanisms to model temporal correlations.
-Graph Convolutional Networks (GCNs) typically construct an adjacency matrix to represent multivariate correlations (Kipf & Welling, 2017; Bruna et al., 2014).
-However, the adjacency matrix is often non-learnable.
-Although the physical locations of sensors are relatively fixed, the correlations among variables are not necessarily constant, which severely affects the performance of GCNs.
-Introducing a learnable adjacency matrix (Wang et al., 2019) can substantially increase computational overhead and lead to model instability (Wu et al., 2020a; Franceschi et al., 2019).
-Therefore, this paper adopts a "weak learning" strategy, whereby multivariate correlations are simply extracted and incorporated as features into tokens to correct temporal features.
-We validate the effectiveness and scalability of the proposed HTME strategy by integrating it into various Transformer variants.
-HTME can be seamlessly combined with various attention mechanisms across a wide spectrum of time series forecasting tasks to achieve optimal performance.
-Our model demonstrates superior performance in short-term sequence forecasting.
-On the selected datasets, PEMS03, PEMS04, and PEMS08, our model achieves the best results compared with the second-best methods.
-In long-term sequence forecasting, our model ranks first or second on eight selected datasets in most cases.
-This demonstrates the effectiveness of the HTME strategy in advancing time series forecasting.
-We also compared the Inference Time, GPU Memory Footprint, and Model Parameters of HTMformer with those of three models that have good forecasting performance: MultiPatchFormer, WPMixer, and PatchTST.
-On the selected datasets, HTMformer achieves superior or equivalent performance with only about 0.2-0.45 times the memory of the three models, and is two to three times faster.
-These results suggest that the HTME model is highly flexible for deployment across a broader spectrum of devices and can better accommodate real-time forecasting tasks, thereby creating greater potential for practical applications and future scalability.
 Figure 8: Examples of forecasts for the ECL dataset with a 96-step predictions.